### v2

# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

Abstract
Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS). Existing lane detection algorithms generally perform well under favorable weather conditions. However, their performance degrades significantly in adverse conditions, such as fog, which increases the risk of traffic accidents. This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments. To address this, we introduce the FoggyLane dataset, captured in real-world foggy scenarios, and synthesize two additional datasets, FoggyCULane and FoggyTusimple, from existing popular lane detection datasets. Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions. Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, with F1-scores of 95.04 on FoggyLane, 79.85 on FoggyCULane, and 96.95 on FoggyTusimple. Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
Index Terms:
ADAS, intelligent vehicle, lane detection, real-time, foggy lane detection dataset, foggy lane detection
I Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving.
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
Figure 1: Lane detection in foggy scenarios for ADAS.[1]
Earlier lane detection algorithms were mainly based on traditional computer vision methods, typically including three steps: image preprocessing, feature extraction, and lane model fitting. Son et al.[2] first detected the vanishing point to adaptively select the ROI, then extracted lane markings based on color features, followed by clustering to find the main lane. Hou et al.[3] used the Canny operator to extract lane edges, filtered out noise with abnormal directions, and applied a progressive probabilistic Hough transform for line detection, filtering out unwanted lines with vanishing point constraints. These feature-based methods are simple, efficient, and suitable for real-time applications, but they also have clear limitations, such as requiring manual parameter tuning and high dependence on clear lane edges. When lane markings are worn, obscured, or subject to environmental changes, accurate detection becomes challenging, and robustness is limited. Improving and optimizing these methods remains necessary to enhance their robustness and adaptability for practical use.
Owing to deep learning’s superior performance in image-based tasks, it now serves as a fundamental approach for many vision tasks. In lane detection, deep learning-based approaches have become mainstream due to their strong feature extraction capabilities and generalization. Compared to traditional methods, deep learning provides higher robustness, better adapting to complex road environments and effectively handling cases with missing or occluded lane markings. LaneNet[4] approaches lane detection as a combination of semantic segmentation and post-processing, first generating a semantic segmentation map of the lanes and then using density clustering to identify each lane instance. However, because of the large number of pixels generated during segmentation and the inability of clustering to leverage GPU acceleration, its real-time performance remains limited. LaneAF[5], building on the semantic segmentation map, introduced a clustering algorithm based on horizontal and vertical vector fields to accelerate post-processing and used DLA[6] as the backbone network to integrate multi-level semantic and spatial information, enhancing feature representation. SCNN[7] leveraged the elongated structure of lane lines by replacing traditional layer-wise convolution with slice convolution, enabling cross-row and cross-column information transfer, which makes it well-suited to line-like lane structures. However, its large parameter size prevents real-time performance. Overall, current networks, whether for semantic or instance segmentation, mostly rely on pixel-by-pixel classification, though lane detection does not necessarily require dense pixel segmentation. To address this, Qin et al.[8] reframed lane detection as a row-based segmentation problem using global features, significantly reducing computational cost and improving detection speed by selecting specific rows and classifying each pixel in the row. This method also leverages the large receptive field of global features, enabling it to handle complex scenarios with missing or occluded lanes.
Apart from segmentation, lane detection also includes detection-based methods, many of which borrow the Anchor-based Methods from object detection[9, 10, 11, 12, 13, 14, 15]. However, due to the elongated shape of lane lines, traditional bounding boxes are not ideal. In lane detection, anchors are often a series of preset starting points and straight lines with different slopes, represented as point sets, with the network predicting the relative distance between the lane line and anchor to optimize line shape and position. PointLaneNet[16] first represents lane lines as a series of points, taking the intersections of evenly spaced horizontal lines as coordinates $\{x_{1},x_{2},\ldots,x_{n}\}$ to describe the lane. Like other YOLO-based methods[9, 10, 17, 18, 19, 20, 21], PointLaneNet divides the image into multiple grids, predicts lane position and classification at each grid center, and removes duplicate predictions through post-processing. However, this grid-based prediction approach has limitations for elongated lanes, reducing detection accuracy. LaneATT[22] addresses this issue by noting the high correlation between lane positions and introducing global features to infer lane position, especially in cases of occlusion or missing lane markers. LaneATT[22] introduced an anchor-based attention mechanism that significantly improves detection speed and accuracy by utilizing global information. CLRNet[23] demonstrated the complementary roles of low- and high-dimensional features in lane detection, proposing Cross Layer Refinement (CLR) to dynamically adjust anchor points through network learning, removing the limitations of fixed anchors to optimize lane detection further. FLAMNet[24] proposes an adaptive line anchor framework integrated with a dual-branch network design to improve the accuracy and adaptability of capturing intricate lane geometries. By leveraging the Decomposed Self-Attention (DSA) mechanism alongside a Horizontal Feature Fusion Module (HFFM), the model enhances the perceptual range of line anchors in lane detection tasks. This approach strengthens robustness against occlusions and incomplete lane markings in challenging environments.
Compared to other detection and segmentation tasks, research on downstream tasks in adverse weather conditions, such as fog, is relatively limited. A conventional solution often applies standard image enhancement techniques, such as classical dehazing methods developed for fog removal and visual quality improvement. Nevertheless, enhanced visual clarity doesn’t always correlate with superior detection outcomes. Some methods combine image enhancement with detection to reduce adverse weather interference. Huang et al.[25] proposed DSNet, incorporating a detection subnet and recovery subnet, where multi-task learning combines dehazing and detection to improve understanding of foggy scenes, resulting in better detection than a standard detection network. Additionally, some methods simulate foggy images from existing datasets to optimize downstream tasks. For example, Nie et al.[26] used the Monodepth2 model and atmospheric scattering model to convert clear lane images from the CULane dataset into foggy lane images, achieving synthetic data augmentation. Testing with SCNN[7] showed that this augmentation strategy improved recognition rates from 76.65% to 86.65% under light fog conditions.
Another approach employs domain adaptation techniques to address the style gap between foggy scenes and clear-weather images[27, 28, 29]. Hnewa et al.[30] assumed that there is domain shift between images captured in normal and adverse weather conditions, proposing a multi-scale domain-adaptive YOLO that supports domain adaptation at different levels in the feature extraction stage. Li et al.[31] proposed a novel transfer learning-based framework for fog-adaptive object detection, which integrates dual-level adaptation (image-level and object-level) to mitigate domain shifts caused by foggy conditions. Notably, the method only requires labeled clear-weather images and unlabeled foggy data for training, eliminating the dependency on annotated foggy datasets.
Traditional dehazing methods, which are based on statistical or physical principles, often suffer from artifacts such as image darkening or brightening. In contrast, end-to-end deep learning methods using Convolutional Neural Networks (CNNs) can capture more complex features but come with higher computational and memory demands. These methods are also highly dependent on diverse training datasets. However, acquiring real paired images under foggy conditions is challenging, and as a result, many existing datasets rely on synthetic images, which can negatively affect algorithm performance in real-world scenarios.
Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS), supporting key functionalities such as lane departure warnings and autonomous navigation. While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents. Furthermore, the lack of specialized datasets tailored to foggy conditions has limited the development of effective deep learning-based solutions for these challenging scenarios.
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. In addition, we generate foggy versions of two widely-used lane detection datasets, CULane [7] and Tusimple [32], creating FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
In response to the unique challenges posed by foggy conditions, we propose a robust real-time lane detection method optimized for such environments. The main contributions of our work are as follows:
  • 1)
    To address the issue of missing global information in foggy images, we design a Global Feature Fusion Module (GFFM) within the backbone network to capture relationships between inputs, improving feature extraction in foggy scenarios.
  • 2)
    To enhance the utilization of structural relationships between lane instances, we introduce a Kernel Feature Fusion Module (KFFM) that learns and predicts the correlations between lane instances, improving lane detection accuracy and robustness.
  • 3)
    To tackle the loss of edge information in foggy images, we incorporate a Low-level Edge Enhancement Module (LEEM) that enhances the model’s attention to edge features, improving the accuracy and stability of lane detection.
  • 4)
    To address the lack of open-source foggy lane detection datasets, we construct FoggyLane, a real-world dataset with 1,423 annotated images, and generate FoggyCULane and FoggyTusimple based on existing datasets. Our method achieves state-of-the-art performance with inference speeds of 192.9 FPS on NVIDIA RTX 3090 and 38.4 FPS on NVIDIA Jetson AGX Orin, confirming its real-time capability and practical applicability in real-world settings.
II Related Work
In recent years, advances in deep learning have substantially improved lane detection performance. Building on these advances, current approaches for lane detection can be categorized into four main paradigms based on their modeling principles: segmentation-based, anchor-based, curve-based, and row-wise-based methods.
II-A Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet[4] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN[7] extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition. RESA[33] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
II-B Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet[34] introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane. LSTR[35] can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information. DBNet[36] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
II-C Anchor-based Lane Detection
Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings. Line-CNN[37] introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors. LaneATT[22] effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module. CLRNet[23] employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features. This hierarchical fusion of semantic information significantly improves detection precision. FLAMNet[24] employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms. This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures. However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
II-D Row-wise-based Lane Detection
Row-wise methods predict lane positions for each image row. E2E-LMD[38] first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions. UFLD[8] utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet[39] utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
III Data Constrution
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [7] and Tusimple [32] datasets to test the generalizability of our methods.
Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.[40]
Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[40]
III-A FoggyLane
In the field of intelligent connected vehicle ADAS systems, existing lane detection algorithms have shown good performance under favorable weather conditions. However, their performance remains significantly inadequate under adverse weather conditions, such as fog. Given that fog is a high-risk scenario for traffic accidents, the limited visibility drastically reduces the driver’s reaction time and decision-making ability, making accidents more likely. However, current benchmark datasets for lane detection, including CULane[7], Tusimple[32], and LLAMAS[41], do not include data from foggy scenarios. To address this gap, we have developed a lane detection dataset, FoggyLane, specifically designed for foggy conditions. Given the randomness of foggy weather, data collection was carried out using a monocular forward-facing camera[42] at different times and locations. This includes the morning of October 26, 2023, in Suzhou, Jiangsu Province, China, and two sessions in Urumqi, Xinjiang Uygur Autonomous Region, China: the morning of December 9, 2023, and the night of January 2, 2024. In these selected representative traffic scenarios, we sampled one frame every frames, resulting in a total of 1,086 RGB images: 444 images (resolution: 1920×1080), 102 images (resolution: 2560×1920), and 540 images (resolution: 2304×1296). To simplify operations and enhance usability during training, we resized all images to a uniform resolution of 1640×590.
However, since the duration of foggy weather conditions is limited, the number of collected foggy weather images is restricted. To improve the model’s performance, we extended our dataset by incorporating additional foggy driving footage. Specifically, we downloaded two foggy driving videos from Bilibili[43], extracting road-driving segments at 20-frame intervals, obtaining 44 and 293 frames, respectively. This added a total of 337 additional RGB images to the dataset. The original resolution of these images was 1920×1080, and they were resized to 1640×590 for consistency.
Figure 4: The different road types of FoggyLane images examples.
As a result, we have constructed FoggyLane, a foggy weather lane detection dataset consisting of 1,423 images. FoggyLane includes six types of road scenes, as shown in Fig. 4, covering Normal, Arrow, Crowd, Curve, Night and Crossroad. We split the dataset into training and testing sets at a ratio of 2:1. To ensure comprehensive learning of scene characteristics, each type of scene is evenly distributed between the training and testing sets to maintain balance.
Meanwhile, We compared FoggyLane with existing lane detection datasets. As shown in Table I, the Caltech-Lanes and Tusimple datasets only cover single traffic scenarios. The LLAMAS [41] dataset focuses on highway scenarios, but its annotations are generated from high-precision maps rather than manually annotated. While the ApolloScape [44] and BDD100k [45] datasets contain multiple traffic scenarios, they lack instance-level annotations, and ApolloScape [44] does not provide publicly available test set annotations. In comparison, although CULane [7] and CurveLanes [46] include a wide range of traffic scenes with comprehensive annotations, they still lack foggy weather scenarios. For lane detection tasks, models trained only on clear weather conditions struggle to handle the challenges posed by fog, such as blurred lane boundaries, faded colors, and reduced image quality. To address these limitations, we developed FoggyLane, a comprehensive foggy weather lane detection dataset that includes 1,423 images featuring various challenging scenarios such as foggy roads during the day, curves, crossroads, and night fog. The dataset is meticulously annotated using the LabelMe tool, providing instance-level annotations, and it is available in Tusimple [32] and CULane [7] formats for broader accessibility.
TABLE I: Comparison Between Our Dataset and Existing Lane Detection Datasets
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [47] 1.2k None 640 x 480 Tusimple [32] 6.4k None 1280 x 720 LLAMAS [41] 100k None 1276 x 717 Appolloscape [44] 110k None 3384 x 2710 BDD100k [45] 100k None 1280 x 720 CULane [7] 133k None 1640 x 590 CurveLanes [46] 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
Our dataset has two key advantages:
  1. 1.
    Pioneering foggy lane detection dataset: FoggyLane is a dataset specifically designed to address foggy weather scenarios. It covers a range of challenging conditions, including foggy congested roads, foggy curves, and foggy night roads. This dataset is crucial for testing the robustness of ADAS under adverse conditions.
  2. 2.
    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[48]. These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
Figure 6: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
III-B FoggyCULane and FoggyTusimple
The CULane[7] dataset is a large-scale and challenging lane detection dataset that is collected from various scenarios, including highways, urban and rural areas in Beijing. Meanwhile, the TuSimple[32] dataset is collected from driving videos on American highways, with most data collected during the day. It encompasses a wide array of road images, capturing varying lighting conditions and diverse traffic scenarios. Both of these datasets are mainstream in the field of lane detection and can comprehensively evaluate the performance of algorithms. Therefore, we have chosen these two datasets to synthesize a foggy dataset, named FoggyCULane and FoggyTuSimple.
The method for generating foggy images is inspired by the model [26], and the complete generation process is shown in Fig. 5. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder[49] is represented as:
$I(x)=J(x)t(x)+A(1-t(x))$ (1)
where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image. $A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
$t(x)=e^{-\beta d(x)}$ (2)
where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter. $d(x)$ represents the distance between the object and the viewer. As this distance increases, the transmittance decreases, resulting in thicker fog.
Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
To calculate the corresponding $d(x)$, we use the S2R-DepthNet[50] that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information. The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image. After passing through the DSA Module and STE Module, as shown in Fig. 5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map. The Depth Prediction (DP) is responsible for depth estimation.
Figure 8: Overview of the proposed network.We introduced a Global Feature Fusion Module(GFFM) into the Swin Transformer[51] backbone to address the lack of global information in foggy images. The neck network employs HRNetV2[52] for multi-scale feature fusion, aggregating features from different scales. The detection head consists of lane instance detection and mask segmentation, using conditional convolution[39] to obtain instance segmentation maps for each lane. Additionally, in order to learn the structural relationships between lanes and address the missing edge information of lanes in foggy scenes, we designed the Kernel Feature Fusion Module (KFFM) and the Low-level Edge Enhancement Module (LEEM).
According to the transmittance formula Eq. (2), we use the dark channel prior[53] to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark }}(x)$ can be defined as:
$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (3)
where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$. In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$. After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$. The original images and the foggy images generated under different $\beta$ values are displayed in Figs. 6 and 7.
IV Methods
IV-A Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods. As illustrated in the Fig. 8, the network comprises three main components: a backbone network, a neck network, and a detection head. To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network. In the neck network, we utilize HRNetV2[52] to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
For the detection head, we employ the concept of dynamic convolution, incorporating both lane instance detection and lane mask segmentation. During instance detection, the network identifies lane instances within the image and predicts a set of dynamic convolution kernel parameters for each lane. These parameters are subsequently applied to the lane mask segmentation head to filter the instance-specific segmentation map for each lane. This approach eliminates the need for complex post-processing steps, enhancing detection speed.
Moreover, to fully exploit the structural relationships between lane lines, we designed a kernel feature fusion module, which allows the network to automatically learn the correlations between predicted lane instances. This contributes to more accurate lane identification and localization. Additionally, to compensate for the loss of edge information in foggy images, we incorporated a low-level edge enhancement module, further improving the network’s ability to detect lane boundaries with precision.
Figure 9: Backbone Network with Global Feature Fusion. The Swin Transformer’s window attention mechanism mainly focuses on local features. Although its layered structure increases the receptive field, it may still fall short in understanding foggy scenes. To address this, we designed a Global Feature Fusion Module (GFFM), replacing the final stage’s window attention with standard multi-head attention to capture global relationships.
IV-B Global Feature Fusion Module
In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision. With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information. As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[54, 55, 56]. However, applying Transformers to the visual domain still faces some challenges. Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities. Additionally, calculating global attention at high resolutions can be computationally intensive.
To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers[51, 57, 58]. Among them, the Swin Transformer[51] combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework. By incorporating a sliding window approach similar to convolution, the Swin Transformer limits self-attention calculations within each sliding window, enabling local information interactions while reducing computational load and improving efficiency. The Swin Transformer also adopts a multi-scale hierarchical structure similar to that of CNNs, dividing the model into multiple stages. In each stage, a Patch Merging layer performs down-sampling to reduce the computational load, followed by the Swin Transformer Block, which applies self-attention within sliding windows. The Swin Transformer Block includes both Window Attention and Shifted Window Attention mechanisms; the former computes self-attention within each window, while the latter enables information exchange across windows.
The Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy. Thus, we chose the Swin Transformer as our backbone network. However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding. Therefore, as illustrated in the Fig. 9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration. Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs. At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
IV-C Kernel Feature Fusion Module
Dynamic convolution is a specialized convolution operation that dynamically adjusts kernel parameters based on the input data. In conventional convolutions, kernel parameters remain fixed after training and are independent of the input data. In contrast, dynamic convolution adapts kernel parameters for each input, enhancing the network’s flexibility and generalization. For specific tasks, dynamic convolution often achieves better performance than standard convolution. CondInst[59] is an instance segmentation framework based on dynamic convolution, which generates convolution kernel parameters dynamically for each instance based on its features and applies them in the segmentation head to filter the instance’s mask. Since each kernel parameter only needs to handle a single instance, this approach reduces the network’s learning burden and convolutional computation load, while also eliminating the ROI operations required in methods like Mask R-CNN, creating a faster, more streamlined instance segmentation framework. Other frameworks, such as SOLO[60], SOLOv2[61], and CondLane[39] , have also applied dynamic convolution in instance segmentation, with CondLane specifically designed for lane instance segmentation, achieving robust performance with its dynamic convolution-based lane detection strategy.
Building on these studies, our lane detection head also incorporates dynamic convolution. This detection head includes two components: lane instance detection and lane mask segmentation. During instance detection, the network identifies lane instances within the image and generates a set of dynamic convolution kernel parameters for each lane. These kernel parameters are applied in the lane mask segmentation head to isolate each lane’s instance segmentation. In typical dynamic convolution-based instance segmentation frameworks, the center position features of each instance are used to generate the dynamic convolution kernels. However, for lane detection, this is less effective, as lanes are elongated structures, making it challenging for center features to capture information along the full lane length. Therefore, following CondLane’s approach, we use the starting position features of each lane to guide the dynamic convolution kernel generation. Specifically, we design an instance detection head in the third stage of the neck network’s feature map, comprising two branches: one branch predicts the starting points of lanes, generating a heatmap for these points; the other branch predicts a set of dynamic convolution kernels for each position, using the kernel at the lane’s starting position for subsequent segmentation.
However, the existing method using lane starting points does not account for the inherent structural and positional relationships among lane instances. In a top-down view, lanes are often parallel, similar in shape, and have consistent spacing. To fully leverage this structural information, we designed a Kernel Feature Fusion Module (KFFM) to enable the network to automatically learn and predict relationships between lane instances. The KFFM is shown in the bottom right of the Fig. 8.
IV-D Low-level Edge Enhanced Module
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking[62]. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 10. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
Figure 10: Comparison of edge information between foggy and normal scenarios.
To address the challenges of lane detection in foggy conditions, we apply global feature fusion within both the backbone network and detection head to enhance high-dimensional global features. Additionally, we use multi-scale fusion to aggregate detail features from low-resolution feature maps. However, this approach alone does not fully compensate for the loss of edge detail inherent in foggy images. Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[7] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[8] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[8] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[22] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[22] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[22] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[35] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[33] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[33] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[23] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[23] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[23] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[23] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[24] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[24] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[36] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[36] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine. * indicates the FPS after acceleration with TensorRT.
Specifically, we add an auxiliary path in the network to learn the edge features of the input image. First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module. To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network. This branch sequentially applies a 3 × 3 and a 1 × 1 convolution. We then compute the cross-entropy loss between the branch output and the edge labels. The detailed structure is illustrated in the bottom-left of the Fig. 8. Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
Figure 11: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[7] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[8] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[8] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[22] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[22] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[22] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[35] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[33] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[33] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[23] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[23] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[23] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[23] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[24] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[24] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[36] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[36] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance. Meanwhile, in the Dazzle Light scenario, our method is only 0.01 percentage points behind the leading performance.
Figure 12: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
V Experiments
V-A Datasets
In the Data Construction chapter, we introduce the FoggyLane dataset, which was constructed from real-world collections for foggy lane detection, along with the FoggyCULane and FoggyTusimple datasets that were synthesized through modeling foggy images. Our method is evaluated on these three datasets.
FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach. It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing. The resolution of all images is also 1640x590 pixels.
FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method. At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases. All visual data maintains a consistent resolution of 1280×720 pixels.
Figure 13: Several metrics of our model’s performance[63] during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
V-B Evaluation Metric
Lane detection tasks are typically regarded as a semantic segmentation problem; therefore, the evaluation metrics for lane detection are often similar to those used in segmentation tasks, specifically the confusion matrix: TP (True Positive), FP (False Positive), FN (False Negative), and TN (True Negative). In the context of lane detection, TP typically represents the correctly predicted lane line pixels, FP denotes incorrectly predicted lane points, FN refers to lane line points that were not detected by the model but are actually present, and TN indicates correctly predicted background pixels. Furthermore, the nature of lane detection tasks introduces some unique characteristics that differentiate their evaluation criteria from those of standard segmentation tasks.
For FoggyLane and FoggyCULane, we adopt the evaluation metrics of SCNN, connecting the labeled or predicted lane points sequentially to form a long line segment with a width of 30 pixels centered around the connecting line. We then calculate the Intersection over Union (IoU) between the predicted area and the ground truth area. If the IoU of the predicted area and the ground truth area exceeds a certain threshold (typically set at 0.5), the lane detection is considered correct; otherwise, it is deemed incorrect. Based on this, we calculate the Precision, Recall, and F1-score, with the following formulas:
$F1=\frac{2\times Precision\times Recall}{Precision+Recall}$ (4)
$Precision=\frac{TP}{TP+FP}$ (5)
$Recall=\frac{TP}{TP+FN}$ (6)
For the FoggyTusimple dataset, we adopt the evaluation metrics of the Tusimple dataset. The official Tusimple dataset provides three standard evaluation metrics: false positive rate (FPR), false negative rate (FNR), and accuracy. The formula for calculating Accuracy is as follows:
$\textit{Accuracy}=\frac{\sum_{\textit{clip}}C_{\textit{clip}}}{\sum_{\textit{clip}}S_{\textit{clip}}}$ (7)
where $C_{\textit{clip}}$ represents the number of correctly predicted lane points in the image, while $S_{\textit{clip}}$ denotes the total number of ground truth lane points. For the pixels in the image, a predicted lane point is considered correct if it falls within 20 pixels to the left or right of the ground truth. Furthermore, if more than 85% of the predicted lane points on a single lane line correspond to the ground truth lane points, the predicted lane is regarded as a true positive.
V-C Implementation Details
We conducted our experiments on the NVIDIA GeForce RTX 3090 platform. To achieve better and more stable performance[64], we first pre-trained the SwinGFFM backbone network on the COCO dataset for 500 epochs. Subsequently, during the training process for lane detection tasks, we loaded the pre-trained weights of the backbone network and resized the input images to 800 × 320. Additionally, we applied various data augmentation techniques to the input images, including rotation, horizontal flipping, brightness adjustment, and random scaling and cropping[65]. During training, the batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay strategy. For detailed network parameters and training settings, please refer to our open-source code.
For objective benchmarking against established lane detection methods including SCNN[7], RESA[33], LaneATT[22], CLRNet[23], FLAMNet[24], LSTR[35], DBNet[36], UFLD[8], and CondLane[39], we replicated these methods on our experimental platform. Given the differences in the architecture of each algorithm, the required batch size, number of training epochs, and learning rate vary, and forcing uniform training settings could negatively impact model performance. Therefore, we adhered to the training configurations provided in the respective papers. Specifically, as the FoggyCULane and FoggyTusimple datasets are extended versions of the CULane and Tusimple datasets, we adopted the original training parameters from CULane and Tusimple when evaluating the algorithms on FoggyCULane and FoggyTusimple. For the FoggyLane dataset, we used the training parameters from CULane and extended the training epochs to ensure sufficient training, thereby obtaining reliable test results.
V-D Quantitative Results
V-D1 Results on FoggyLane
We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods. As shown in the Tabel II, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
From the Table II, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
V-D2 Results on FoggyCULane
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
V-D3 Results on FoggyTusimple
As shown in the Table IV, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods. Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%. This further demonstrates the effectiveness of our method in foggy conditions.
TABLE IV: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[7] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[8] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[8] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[22] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[22] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[22] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[35] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[33] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[33] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[23] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[23] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[23] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[24] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[24] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[36] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[36] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
V-E Ablation Study
To explore the effectiveness of each part of the model, we conducted ablation studies on both real and synthetic datasets, namely FoggyLane and FoggyCULane. We have studied the roles of the GFFM, KFFM, and LEEM modules in the model, with Swin-t as the backbone network. As shown in the Tables. V and VI, adding each module boosts the model’s performance on both datasets. Among them, GFFM contributes the most to the model’s improvement, with increases of 1.07 on FoggyLane and 0.41 on FoggyCULane. This indicates that GFFM can effectively capture both local and global information, enriching the feature representation. Building on the foundation of GFFM, the addition of the KFFM module further increases the F1-score by 0.63 on FoggyLane and by 0.26 on FoggyCULane. GFFM is capable of automatically learning the associations between lane instance instances, making full use of relevant information. On this basis, the introduction of the LEEM module enhances the network’s focus on edge information. The F1-score increases by 0.28 on FoggyLane and by 0.14 on FoggyCULane.
TABLE V: Ablation experiments of different components on the FoggyLane dataset.
  Model GFFM KFFM LEEM F1-score Gflops(G)
Baseline 90.36 13.77
91.43 11.94
92.06 11.97
Ours 92.34$\uparrow$ 12.01
TABLE VI: Ablation experiments of different components on the FoggyCULane dataset.
  Model GFFM KFFM LEEM F1-score Gflops(G)
Baseline 77.85 13.77
78.26 11.94
78.52 11.97
Ours 78.66$\uparrow$ 12.01
V-F Model deployment
Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin. After experimental validation, it has been demonstrated that our method can run accurately on the vehicle’s edge computing platform with limited computational resources, achieving a performance of 38.4 FPS.
To further evaluate the performance of our algorithm in real-world scenarios, in addition to experiments on existing datasets, we also deployed and tested the model on the vehicle. Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig. 14. This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit. The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
VI Conclusion
This study focuses on lane detection in foggy conditions, aiming to address the range of challenges posed by foggy environments, which is crucial for ensuring driving safety in fog and enhancing Advanced Driver Assistance System (ADAS) capabilities. To address the current lack of publicly available datasets for lane detection in foggy conditions, we constructed a real-world foggy lane dataset, FoggyLane, and synthesized two additional foggy datasets, FoggyCULane and FoggyTusimple, based on the popular CULane and Tusimple datasets.
In response to the challenges of lane detection in foggy scenarios, we propose an end-to-end fog-enhanced lane detection method. First, this method incorporates a Global Feature Fusion Module in the backbone network to address the absence of global spatial information in foggy scenes. Second, we employ a dynamic convolution-based detection head for instance-sensitive lane segmentation, integrating a Kernel Feature Fusion Module within the detection head to automatically capture associations between lane instances. Additionally, a Low-Level Edge Enhancement Module was added to increase the network’s focus on edge information, mitigating the loss of edge details and improving lane detection accuracy under foggy conditions.
Comparative experiments on the FoggyLane, FoggyCULane, and FoggyTusimple datasets with other advanced lane detection methods demonstrate that our approach achieves superior performance in both clear and foggy conditions, significantly improving lane detection accuracy. Furthermore, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, further validating its practical applicability in real-world scenarios.
Acknowledgments
This project is jointly supported by National Natural Science Foundation of China (Nos. 52172350, W2421069 and 51775565), the Guangdong Basic and Applied Research Foundation (No. 2022B1515120072), the Guangzhou Science and Technology Plan Project (No. 2024B01W0079), the Nansha Key RD Program (No. 2022ZD014), the Science and Technology Planning Project of Guangdong Province (No. 2023B1212060029).We would like to release our source code and dataset.

### v3

@@
-A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions
+
 Abstract
@@
 Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
-Index Terms:
-ADAS, intelligent vehicle, lane detection, real-time, foggy lane detection dataset, foggy lane detection
-I Introduction
+keywords:
+[…]
+1 Introduction
 Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents.
@@
 Under good weather and clear lane markings, current lane detection algorithms perform well.
-However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
-In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences.
+However, in complex road environments[1] and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
+In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences[2].
 Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments.
@@
 Improving and optimizing these methods remains necessary to enhance their robustness and adaptability for practical use.
-Owing to deep learning’s superior performance in image-based tasks, it now serves as a fundamental approach for many vision tasks.
-In lane detection, deep learning-based approaches have become mainstream due to their strong feature extraction capabilities and generalization.
+Owing to deep learning’s superior performance in image-based tasks, it now serves as a fundamental approach for many vision tasks[6, 7].
+In lane detection, deep learning-based approaches[8] have become mainstream due to their strong feature extraction capabilities and generalization.
 Compared to traditional methods, deep learning provides higher robustness, better adapting to complex road environments and effectively handling cases with missing or occluded lane markings.
@@
 This method also leverages the large receptive field of global features, enabling it to handle complex scenarios with missing or occluded lanes.
-Apart from segmentation, lane detection also includes detection-based methods, many of which borrow the Anchor-based Methods from object detection[9, 10, 11, 12, 13, 14, 15].
+Apart from segmentation, lane detection also includes detection-based methods, many of which borrow the Anchor-based Methods from object detection[14, 15, 16, 17].
 However, due to the elongated shape of lane lines, traditional bounding boxes are not ideal.
@@
 PointLaneNet[18] first represents lane lines as a series of points, taking the intersections of evenly spaced horizontal lines as coordinates $\{x_{1},x_{2},\ldots,x_{n}\}$ to describe the lane.
-Like other YOLO-based methods[9, 10, 17, 18, 19, 20, 21], PointLaneNet divides the image into multiple grids, predicts lane position and classification at each grid center, and removes duplicate predictions through post-processing.
+Like other YOLO-based methods[14, 15], PointLaneNet divides the image into multiple grids, predicts lane position and classification at each grid center, and removes duplicate predictions through post-processing.
 However, this grid-based prediction approach has limitations for elongated lanes, reducing detection accuracy.
@@
 Testing with SCNN[12] showed that this augmentation strategy improved recognition rates from 76.65% to 86.65% under light fog conditions.
-Another approach employs domain adaptation techniques to address the style gap between foggy scenes and clear-weather images[27, 28, 29].
+Another approach employs domain adaptation techniques to address the style gap between foggy scenes and clear-weather images[24, 25].
 Hnewa et al.[26] assumed that there is domain shift between images captured in normal and adverse weather conditions, proposing a multi-scale domain-adaptive YOLO that supports domain adaptation at different levels in the feature extraction stage.
@@
 Our method achieves state-of-the-art performance with inference speeds of 192.9 FPS on NVIDIA RTX 3090 and 38.4 FPS on NVIDIA Jetson AGX Orin, confirming its real-time capability and practical applicability in real-world settings.
-II Related Work
+2 Related Work
 In recent years, advances in deep learning have substantially improved lane detection performance.
 Building on these advances, current approaches for lane detection can be categorized into four main paradigms based on their modeling principles: segmentation-based, anchor-based, curve-based, and row-wise-based methods.
-II-A Segmentation-based Lane Detection
+2.1 Segmentation-based Lane Detection
 Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem.
@@
 Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
-II-B Curve-based Lane Detection
+2.2 Curve-based Lane Detection
 Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations.
@@
 However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
-II-C Anchor-based Lane Detection
+2.3 Anchor-based Lane Detection
 Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings.
@@
 However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
-II-D Row-wise-based Lane Detection
+2.4 Row-wise-based Lane Detection
 Row-wise methods predict lane positions for each image row.
@@
 The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
-III Data Constrution
-Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models.
+3 Data Constrution
+Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models[36].
 In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions.
 Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging.
-Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability.
+Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability[37, 38].
 However, there is currently no publicly available foggy weather lane detection dataset.
@@
 Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[39]
-III-A FoggyLane
+3.1 FoggyLane
 In the field of intelligent connected vehicle ADAS systems, existing lane detection algorithms have shown good performance under favorable weather conditions.
@@
 To address this gap, we have developed a lane detection dataset, FoggyLane, specifically designed for foggy conditions.
-Given the randomness of foggy weather, data collection was carried out using a monocular forward-facing camera[42] at different times and locations.
+Given the randomness of foggy weather, data collection was carried out using a monocular forward-facing camera at different times and locations.
 This includes the morning of October 26, 2023, in Suzhou, Jiangsu Province, China, and two sessions in Urumqi, Xinjiang Uygur Autonomous Region, China: the morning of December 9, 2023, and the night of January 2, 2024.
-In these selected representative traffic scenarios, we sampled one frame every frames, resulting in a total of 1,086 RGB images: 444 images (resolution: 1920×1080), 102 images (resolution: 2560×1920), and 540 images (resolution: 2304×1296).
-To simplify operations and enhance usability during training, we resized all images to a uniform resolution of 1640×590.
+In these selected representative traffic scenarios, we sampled one frame every frames, resulting in a total of 1,086 RGB images: 444 images (resolution: $1920\times 1080$), 102 images (resolution: $2560\times 1920$), and 540 images (resolution: $2304\times 1296$).
+To simplify operations and enhance usability during training, we resized all images to a uniform resolution of $1640\times 590$.
 However, since the duration of foggy weather conditions is limited, the number of collected foggy weather images is restricted.
@@
 This added a total of 337 additional RGB images to the dataset.
-The original resolution of these images was 1920×1080, and they were resized to 1640×590 for consistency.
+The original resolution of these images was $1920\times 1080$, and they were resized to $1640\times 590$ for consistency.
 Figure 4: The different road types of FoggyLane images examples.
@@
 Meanwhile, We compared FoggyLane with existing lane detection datasets.
-As shown in Table I, the Caltech-Lanes and Tusimple datasets only cover single traffic scenarios.
+As shown in Table 1, the Caltech-Lanes and Tusimple datasets only cover single traffic scenarios.
 The LLAMAS [40] dataset focuses on highway scenarios, but its annotations are generated from high-precision maps rather than manually annotated.
@@
 The dataset is meticulously annotated using the LabelMe tool, providing instance-level annotations, and it is available in Tusimple [28] and CULane [12] formats for broader accessibility.
-TABLE I: Comparison Between Our Dataset and Existing Lane Detection Datasets
+Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
 Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [45] 1.2k None 640 x 480 Tusimple [28] 6.4k None 1280 x 720 LLAMAS [40] 100k None 1276 x 717 Appolloscape [42] 110k None 3384 x 2710 BDD100k [43] 100k None 1280 x 720 CULane [12] 133k None 1640 x 590 CurveLanes [44] 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
 Our dataset has two key advantages:
-  1. 1.
+  • 1)
     Pioneering foggy lane detection dataset: FoggyLane is a dataset specifically designed to address foggy weather scenarios.
@@
 This dataset is crucial for testing the robustness of ADAS under adverse conditions.
-  2. 2.
+  • 2)
     Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[46].
@@
 Figure 6: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
-III-B FoggyCULane and FoggyTusimple
+3.2 FoggyCULane and FoggyTusimple
 The CULane[12] dataset is a large-scale and challenging lane detection dataset that is collected from various scenarios, including highways, urban and rural areas in Beijing.
@@
 6 and 7.
-IV Methods
-IV-A Overall structure
+4 Methods
+4.1 Overall structure
 To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods.
@@
 To address this, we designed a Global Feature Fusion Module (GFFM), replacing the final stage’s window attention with standard multi-head attention to capture global relationships.
-IV-B Global Feature Fusion Module
+4.2 Global Feature Fusion Module
 In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision.
-With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information.
-As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[54, 55, 56].
+With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information[52].
+As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[53, 54, 55, 56].
 However, applying Transformers to the visual domain still faces some challenges.
@@
 At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
-IV-C Kernel Feature Fusion Module
+4.3 Kernel Feature Fusion Module
 Dynamic convolution is a specialized convolution operation that dynamically adjusts kernel parameters based on the input data.
@@
 8.
-IV-D Low-level Edge Enhanced Module
+4.4 Low-level Edge Enhanced Module
 Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast.
 This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately.
-Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking[62].
+Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking.
 Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig.
@@
 Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
-TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
+Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
   Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[12] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[13] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[13] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[19] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[19] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[19] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[31] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[29] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[29] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[35] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[35] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[35] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[20] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[20] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[20] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[20] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[21] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[21] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[32] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[32] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
@@
 To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network.
-This branch sequentially applies a 3 × 3 and a 1 × 1 convolution.
+This branch sequentially applies a $3\times 3$ and a $1\times 1$ convolution.
 We then compute the cross-entropy loss between the branch output and the edge labels.
@@
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
+Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
   Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[12] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[13] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[13] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[19] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[19] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[19] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[31] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[29] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[29] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[35] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[35] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[35] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[20] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[20] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[20] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[20] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[21] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[21] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[32] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[32] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
@@
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-V Experiments
-V-A Datasets
+5 Experiments
+5.1 Datasets
 In the Data Construction chapter, we introduce the FoggyLane dataset, which was constructed from real-world collections for foggy lane detection, along with the FoggyCULane and FoggyTusimple datasets that were synthesized through modeling foggy images.
 Our method is evaluated on these three datasets.
-FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
+FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing.
+All images are of resolution $1640\times 590$ pixels.
 FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach.
 It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing.
-The resolution of all images is also 1640x590 pixels.
+The resolution of all images is also $1640\times 590$ pixels.
 FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method.
 At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases.
-All visual data maintains a consistent resolution of 1280×720 pixels.
-Figure 13: Several metrics of our model’s performance[63] during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
-V-B Evaluation Metric
+All visual data maintains a consistent resolution of $1280\times 720$ pixels.
+Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
+5.2 Evaluation Metric
 Lane detection tasks are typically regarded as a semantic segmentation problem; therefore, the evaluation metrics for lane detection are often similar to those used in segmentation tasks, specifically the confusion matrix: TP (True Positive), FP (False Positive), FN (False Negative), and TN (True Negative).
@@
 Furthermore, the nature of lane detection tasks introduces some unique characteristics that differentiate their evaluation criteria from those of standard segmentation tasks.
-For FoggyLane and FoggyCULane, we adopt the evaluation metrics of SCNN, connecting the labeled or predicted lane points sequentially to form a long line segment with a width of 30 pixels centered around the connecting line.
+For FoggyLane and FoggyCULane, we adopt the evaluation metrics of SCNN[12], connecting the labeled or predicted lane points sequentially to form a long line segment with a width of 30 pixels centered around the connecting line.
 We then calculate the Intersection over Union (IoU) between the predicted area and the ground truth area.
@@
 Furthermore, if more than 85% of the predicted lane points on a single lane line correspond to the ground truth lane points, the predicted lane is regarded as a true positive.
-V-C Implementation Details
+5.3 Implementation Details
 We conducted our experiments on the NVIDIA GeForce RTX 3090 platform.
@@
 Subsequently, during the training process for lane detection tasks, we loaded the pre-trained weights of the backbone network and resized the input images to 800 × 320.
-Additionally, we applied various data augmentation techniques to the input images, including rotation, horizontal flipping, brightness adjustment, and random scaling and cropping[65].
+Additionally, we applied various data augmentation techniques to the input images, including rotation, horizontal flipping, brightness adjustment, and random scaling and cropping.
 During training, the batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay strategy.
 For detailed network parameters and training settings, please refer to our open-source code.
-For objective benchmarking against established lane detection methods including SCNN[7], RESA[33], LaneATT[22], CLRNet[23], FLAMNet[24], LSTR[35], DBNet[36], UFLD[8], and CondLane[39], we replicated these methods on our experimental platform.
+For objective benchmarking against established lane detection methods including SCNN, RESA, LaneATT, CLRNet, FLAMNet, LSTR, DBNet, UFLD, and CondLane, we replicated these methods on our experimental platform.
 Given the differences in the architecture of each algorithm, the required batch size, number of training epochs, and learning rate vary, and forcing uniform training settings could negatively impact model performance.
@@
 For the FoggyLane dataset, we used the training parameters from CULane and extended the training epochs to ensure sufficient training, thereby obtaining reliable test results.
-V-D Quantitative Results
-V-D1 Results on FoggyLane
+5.4 Quantitative Results
+5.4.1 Results on FoggyLane
 We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods.
-As shown in the Tabel II, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
-From the Table II, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively.
+As shown in the Tabel 2, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
+From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively.
 These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%.
@@
 Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
-V-D2 Results on FoggyCULane
+5.4.2 Results on FoggyCULane
 We compared our method with other advanced lane detection methods on the FoggyCULane dataset.
-As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
+As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
 Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
-V-D3 Results on FoggyTusimple
-As shown in the Table IV, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods.
+5.4.3 Results on FoggyTusimple
+As shown in the Table 4, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods.
 Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%.
 This further demonstrates the effectiveness of our method in foggy conditions.
-TABLE IV: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
+Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
     Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[12] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[13] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[13] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[19] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[19] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[19] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[31] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[29] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[29] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[35] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[35] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[35] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[20] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[20] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[20] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[21] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[21] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[32] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[32] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
-V-E Ablation Study
+5.5 Ablation Study
 To explore the effectiveness of each part of the model, we conducted ablation studies on both real and synthetic datasets, namely FoggyLane and FoggyCULane.
@@
 As shown in the Tables.
-V and VI, adding each module boosts the model’s performance on both datasets.
+5 and 6, adding each module boosts the model’s performance on both datasets.
 Among them, GFFM contributes the most to the model’s improvement, with increases of 1.07 on FoggyLane and 0.41 on FoggyCULane.
@@
 The F1-score increases by 0.28 on FoggyLane and by 0.14 on FoggyCULane.
-TABLE V: Ablation experiments of different components on the FoggyLane dataset.
+Table 5: Ablation experiments of different components on the FoggyLane dataset.
   Model GFFM KFFM LEEM F1-score Gflops(G)
@@
 Ours 92.34$\uparrow$ 12.01
-TABLE VI: Ablation experiments of different components on the FoggyCULane dataset.
+Table 6: Ablation experiments of different components on the FoggyCULane dataset.
   Model GFFM KFFM LEEM F1-score Gflops(G)
@@
 Ours 78.66$\uparrow$ 12.01
-V-F Model deployment
+5.6 Model deployment
 Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin.
@@
 The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
-VI Conclusion
+6 Conclusion
 This study focuses on lane detection in foggy conditions, aiming to address the range of challenges posed by foggy environments, which is crucial for ensuring driving safety in fog and enhancing Advanced Driver Assistance System (ADAS) capabilities.
@@
 2023B1212060029).We would like to release our source code and dataset.
+Compliance with ethics guidelines
+Ronghui Zhang, Yuhang Ma, Tengfei Li, Ziyu Lin, Yueying Wu, Junzhou Chen, Lin Zhang, Jia Hu,Tony Z.
+Qiu and Konghui Guo declare that they have no conflict of interest or financial conflicts to disclose.

### v4

@@
 Under good weather and clear lane markings, current lane detection algorithms perform well.
-However, in complex road environments[1] and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
+However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
 In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences[1].
@@
 Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud.
-This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
+This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions such as accelerating, braking, or steering[2, 3] to ensure safety in challenging environments like fog.
+Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2).
+In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments.
+Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions[4].
+Therefore, this research not only focuses on end-to-end autonomous driving technologies but also aims to improve lane detection performance in complex environments, providing stronger safety support and driving assistance for L0, L1, and L2 vehicles in adverse conditions.
 Figure 1: Lane detection in foggy scenarios for ADAS.[5]
@@
 Improving and optimizing these methods remains necessary to enhance their robustness and adaptability for practical use.
-Owing to deep learning’s superior performance in image-based tasks, it now serves as a fundamental approach for many vision tasks[6, 7].
-In lane detection, deep learning-based approaches[8] have become mainstream due to their strong feature extraction capabilities and generalization.
+Owing to the rapid development of deep learning, it has been widely applied in various autonomous driving tasks.
+For example, in traffic object detection, Wang and Zhang[8] proposed CRRFNet, an adaptive method that combines camera and radar fusion, improving detection accuracy and robustness in noisy or abnormal scenes.
+In accident detection and classification, Yang et al.[9] developed a deep learning model that utilizes multi-vehicle trajectory data for detecting and classifying accidents on freeways, achieving up to 100% detection accuracy.
+For 3D object detection under adverse weather, Qi et al.[10] introduced a novel geometric information constraint network (GIC-Net), enhancing the performance of 3D object detection by incorporating geometric location and line feature constraints, demonstrating superior results under challenging weather conditions.
+In addition to these applications, lane detection has also seen significant advancements, with deep learning-based approaches becoming mainstream due to their strong feature extraction capabilities and generalization.
 Compared to traditional methods, deep learning provides higher robustness, better adapting to complex road environments and effectively handling cases with missing or occluded lane markings.
@@
 This approach strengthens robustness against occlusions and incomplete lane markings in challenging environments.
+In addition to these methods, trajectory data has also been explored for lane identification.
+Zhao and Roncoli[24] proposed a constrained spectral clustering method to identify lane information from high-resolution trajectory data, significantly improving results even in the presence of varying lane numbers or lane-changing maneuvers.
+This approach demonstrates the potential of leveraging trajectory data for lane detection, especially in more dynamic environments.
 Compared to other detection and segmentation tasks, research on downstream tasks in adverse weather conditions, such as fog, is relatively limited.
@@
 3 Data Constrution
-Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models[36].
+Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models.
 In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions.
 Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging.
-Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability[37, 38].
+Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability.
 However, there is currently no publicly available foggy weather lane detection dataset.
@@
 However, since the duration of foggy weather conditions is limited, the number of collected foggy weather images is restricted.
-To improve the model’s performance, we extended our dataset by incorporating additional foggy driving footage.
-Specifically, we downloaded two foggy driving videos from Bilibili[41], extracting road-driving segments at 20-frame intervals, obtaining 44 and 293 frames, respectively.
+To improve the model’s performance, we extended our dataset by incorporating additional foggy driving scenes.
+Specifically, we downloaded two foggy driving videos from YouTube[41], extracting road-driving segments at 20-frame intervals, obtaining 44 and 293 frames, respectively.
 This added a total of 337 additional RGB images to the dataset.
 The original resolution of these images was $1920\times 1080$, and they were resized to $1640\times 590$ for consistency.
-Figure 4: The different road types of FoggyLane images examples.
 As a result, we have constructed FoggyLane, a foggy weather lane detection dataset consisting of 1,423 images.
@@
 To ensure comprehensive learning of scene characteristics, each type of scene is evenly distributed between the training and testing sets to maintain balance.
+Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
+Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [42] 1.2k None 640 x 480 Tusimple [31] 6.4k None 1280 x 720 LLAMAS [40] 100k None 1276 x 717 Appolloscape [43] 110k None 3384 x 2710 BDD100k [44] 100k None 1280 x 720 CULane [14] 133k None 1640 x 590 CurveLanes [45] 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
+Figure 4: The different road types of FoggyLane images examples.
 Meanwhile, We compared FoggyLane with existing lane detection datasets.
@@
 The dataset is meticulously annotated using the LabelMe tool, providing instance-level annotations, and it is available in Tusimple [31] and CULane [14] formats for broader accessibility.
-Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
-Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [45] 1.2k None 640 x 480 Tusimple [28] 6.4k None 1280 x 720 LLAMAS [40] 100k None 1276 x 717 Appolloscape [42] 110k None 3384 x 2710 BDD100k [43] 100k None 1280 x 720 CULane [12] 133k None 1640 x 590 CurveLanes [44] 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
 Our dataset has two key advantages:
@@
 3.2 FoggyCULane and FoggyTusimple
-The CULane[12] dataset is a large-scale and challenging lane detection dataset that is collected from various scenarios, including highways, urban and rural areas in Beijing.
-Meanwhile, the TuSimple[28] dataset is collected from driving videos on American highways, with most data collected during the day.
-It encompasses a wide array of road images, capturing varying lighting conditions and diverse traffic scenarios.
-Both of these datasets are mainstream in the field of lane detection and can comprehensively evaluate the performance of algorithms.
-Therefore, we have chosen these two datasets to synthesize a foggy dataset, named FoggyCULane and FoggyTuSimple.
+Currently, numerous datasets are available for lane detection.
+Among them, we select two widely adopted and representative benchmarks: CULane[14] and TuSimple[31].
+The CULane dataset [14] is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing.
+In contrast, the TuSimple dataset [31] primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
+Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
 The method for generating foggy images is inspired by the model [26], and the complete generation process is shown in Fig.
@@
 The Depth Prediction (DP) is responsible for depth estimation.
-Figure 8: Overview of the proposed network.We introduced a Global Feature Fusion Module(GFFM) into the Swin Transformer[49] backbone to address the lack of global information in foggy images.
-The neck network employs HRNetV2[50] for multi-scale feature fusion, aggregating features from different scales.
-The detection head consists of lane instance detection and mask segmentation, using conditional convolution[35] to obtain instance segmentation maps for each lane.
-Additionally, in order to learn the structural relationships between lanes and address the missing edge information of lanes in foggy scenes, we designed the Kernel Feature Fusion Module (KFFM) and the Low-level Edge Enhancement Module (LEEM).
+Figure 8: Overview of the proposed network.
 According to the transmittance formula Eq.
@@
 Figure 9: Backbone Network with Global Feature Fusion.
-The Swin Transformer’s window attention mechanism mainly focuses on local features.
-Although its layered structure increases the receptive field, it may still fall short in understanding foggy scenes.
-To address this, we designed a Global Feature Fusion Module (GFFM), replacing the final stage’s window attention with standard multi-head attention to capture global relationships.
 4.2 Global Feature Fusion Module
 In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision.
-With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information[52].
-As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[53, 54, 55, 56].
+With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information.
+As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[51, 52, 53].
 However, applying Transformers to the visual domain still faces some challenges.
@@
 Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
-Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
-  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[12] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[13] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[13] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[19] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[19] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[19] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[31] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[29] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[29] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[35] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[35] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[35] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[20] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[20] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[20] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[20] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[21] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[21] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[32] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[32] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
-  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
-* indicates the FPS after acceleration with TensorRT.
 Specifically, we add an auxiliary path in the network to learn the edge features of the input image.
@@
 Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
-Figure 11: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyLane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
-  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[12] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[13] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[13] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[19] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[19] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[19] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[31] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[29] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[29] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[35] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[35] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[35] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[20] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[20] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[20] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[20] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[21] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[21] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[32] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[32] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
-  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
-Meanwhile, in the Dazzle Light scenario, our method is only 0.01 percentage points behind the leading performance.
-Figure 12: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyCULane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+4.5 Loss functions
+Based on the above, our method involves a total of five prediction tasks: lane start point prediction, row-based segmentation prediction, vertical range prediction, offset prediction, and edge prediction.
+The loss functions for each task are detailed below.
+4.5.1 Lane start point prediction
+In feature maps with a large receptive field and low resolution, the lane line start point is difficult to be accurately represented by a single pixel location.
+Moreover, in such large receptive field feature maps, pixels near the target point exhibit similar features.
+If these pixels are directly labeled as negative samples, it will interfere with the network training and make it difficult for the network to learn and optimize.
+Therefore, this method uses a Gaussian kernel function to generate a heatmap label for the lane line start point.
+Each element in the heatmap label represents the confidence that the position corresponds to the target point; the closer to the target point, the higher the confidence, and the farther away, the lower the confidence.
+In other words, the ground truth is softly labeled with a Gaussian function, making it easier for the network to converge.
+Suppose $(x,y)$ denotes a position on the low-resolution feature map, and $(p_{x},p_{y})$ represents the coordinate of the target point on the low-resolution feature map.
+Then, the ground truth heatmap $GT$ value at position $(x,y)$ is defined as:
+$GT(x,y)=\exp\left(-\frac{(x-p_{x})^{2}+(y-p_{y})^{2}}{2\sigma^{2}}\right)$ (4)
+where $\sigma$ controls the decay rate of the Gaussian kernel. A larger $\sigma$ results in slower confidence decay and thus a larger heatmap area.
+This method combines the Gaussian heatmaps of all lane line start points within the image into a single heatmap.
+When multiple heatmaps respond at the same location, the maximum value is retained to form the final ground truth heatmap.
+Then, the Focal Loss is used to calculate the loss between the predicted heatmap and the ground truth heatmap.
+Let the predicted heatmap value at position $(x,y)$ be $P_{xy}$, and let $N$ be the total number of pixels in the heatmap.
+The Focal Loss between the predicted heatmap and the ground truth heatmap is defined as:
+$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$ (5)
+4.5.2 Row-based segmentation prediction
+Row-based segmentation predicts a horizontal coordinate $\mathrm{E}\left(\hat{x}_{i}\right)$ for each row.
+For this task, our method uses the L1 loss function to optimize the predicted coordinates.
+Suppose $V$ denotes the effective vertical range of the lane line, $N_{v}$ is the number of effective rows, and $x_{i}$ is the ground truth horizontal coordinate.
+The loss function is defined as:
+$L_{\text{row}}=\frac{1}{N_{v}}\sum_{i\in V}\left|\mathrm{E}\left(\hat{x}_{i}\right)-x_{i}\right|$ (6)
+4.5.3 Vertical range prediction
+For the label of the lane line vertical range vector prediction, our method uses a $1\times Y$ vector representation.
+If the lane line passes through the $Y_{i}$ row, its value is 1; otherwise, it is 0.
+Thus, this can be regarded as a segmentation task.
+Our method optimizes it using the cross-entropy loss function.
+Suppose $v_{i}$ denotes the predicted probability that the $i$-th row is positive, then the loss function is defined as:
+$L_{\text{range}}=\sum\left(-y_{gt}^{i}\log(v_{i})-(1-y_{gt}^{i})\log(1-v_{i})\right)$ (7)
+4.5.4 Offset prediction
+Firstly, the generation of ground truth labels for offset prediction is required.
+For each lane line, the proposed method constructs the ground truth offset as the horizontal displacement relative to the lane centerline for each grid cell within a specified width.
+The predicted offsets are optimized using an L1 loss function.
+Let $\delta_{xy}$ denote the ground truth offset at coordinate $(x,y)$, $\hat{\delta}_{xy}$ the predicted offset, and $\Omega$ the region surrounding the lane line.
+The loss function is then formulated as follows:
+$L_{\text{offset}}=\frac{1}{N_{\Omega}}\sum_{(x,y)\in\Omega}\left|\hat{\delta}_{xy}-\delta_{xy}\right|$ (8)
+4.5.5 Edge prediction
+For the task of edge enhancement, the method treats it as an edge segmentation problem, which is essentially a binary classification task.
+The labels at edge locations are assigned a value of 1, while non-edge locations are assigned 0.
+Accordingly, the cross-entropy loss function is employed for optimization.
+Let $e_{gt}^{i}$ denote the label of pixel $i$, and $p_{i}$ represent the predicted probability of the pixel being positive.
+The loss function is defined as:
+$L_{\text{edge}}=\sum_{i}\left(-e_{gt}^{i}\log(p_{i})-(1-e_{gt}^{i})\log(1-p_{i})\right)$ (9)
+4.5.6 Total loss
+To balance the influence of different tasks on the overall training of the network, a weighted fusion of the loss functions for each sub-task is employed.
+The total loss function is formulated as follows:
+$L_{\text{total}}=\alpha L_{hm}+\beta L_{\text{row}}+\gamma L_{\text{range}}+\delta L_{\text{offset}}+\varepsilon L_{\text{edge}}$ (10)
+where the network achieves optimal performance when the weights are set as $\alpha=1$, $\beta=1$, $\gamma=1$, $\delta=0.4$, and $\varepsilon=0.1$.
 5 Experiments
@@
 All visual data maintains a consistent resolution of $1280\times 720$ pixels.
-Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
 5.2 Evaluation Metric
@@
 For detailed network parameters and training settings, please refer to our open-source code.
+Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
+  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[14] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[15] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[15] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[21] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[21] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[21] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[32] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[32] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[38] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[38] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[38] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[22] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[22] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[22] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[22] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[23] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[23] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
+  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
+* indicates the FPS after acceleration with TensorRT.
+Figure 11: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyLane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
+  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[14] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[15] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[15] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[21] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[21] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[21] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[32] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[32] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[38] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[38] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[38] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[22] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[22] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[22] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[22] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[23] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[23] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
+  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
+Meanwhile, in the Dazzle Light scenario, our method is only 0.01 percentage points behind the leading performance.
+Figure 12: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyCULane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
 For objective benchmarking against established lane detection methods including SCNN, RESA, LaneATT, CLRNet, FLAMNet, LSTR, DBNet, UFLD, and CondLane, we replicated these methods on our experimental platform.
@@
 Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
+Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
 5.4.3 Results on FoggyTusimple
@@
 After experimental validation, it has been demonstrated that our method can run accurately on the vehicle’s edge computing platform with limited computational resources, achieving a performance of 38.4 FPS.
-To further evaluate the performance of our algorithm in real-world scenarios, in addition to experiments on existing datasets, we also deployed and tested the model on the vehicle.
+After considering the deployment of several edge computing tasks[61, 62], we further evaluate the performance of our algorithm in real-world scenarios.
+In addition to experiments on existing datasets, we also deployed and tested the model on the vehicle.
 Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig.

### v5

@@
-
+A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions
 Abstract
@@
 Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
-keywords:
-[…]
-1 Introduction
+Index Terms:
+ADAS, intelligent vehicle, foggy lane detection dataset, foggy lane detection, global attention, edge enhancement
+I Introduction
 Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents.
-As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving.
+As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving[1].
 Under good weather and clear lane markings, current lane detection algorithms perform well.
 However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
-In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences[1].
+In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences.
 Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments.
@@
 Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud.
-This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions such as accelerating, braking, or steering[2, 3] to ensure safety in challenging environments like fog.
-Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2).
-In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments.
-Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions[4].
-Therefore, this research not only focuses on end-to-end autonomous driving technologies but also aims to improve lane detection performance in complex environments, providing stronger safety support and driving assistance for L0, L1, and L2 vehicles in adverse conditions.
+This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
 Figure 1: Lane detection in foggy scenarios for ADAS.[2]
-Earlier lane detection algorithms were mainly based on traditional computer vision methods, typically including three steps: image preprocessing, feature extraction, and lane model fitting.
-Son et al.[6] first detected the vanishing point to adaptively select the ROI, then extracted lane markings based on color features, followed by clustering to find the main lane.
-Hou et al.[7] used the Canny operator to extract lane edges, filtered out noise with abnormal directions, and applied a progressive probabilistic Hough transform for line detection, filtering out unwanted lines with vanishing point constraints.
-These feature-based methods are simple, efficient, and suitable for real-time applications, but they also have clear limitations, such as requiring manual parameter tuning and high dependence on clear lane edges.
-When lane markings are worn, obscured, or subject to environmental changes, accurate detection becomes challenging, and robustness is limited.
-Improving and optimizing these methods remains necessary to enhance their robustness and adaptability for practical use.
-Owing to the rapid development of deep learning, it has been widely applied in various autonomous driving tasks.
-For example, in traffic object detection, Wang and Zhang[8] proposed CRRFNet, an adaptive method that combines camera and radar fusion, improving detection accuracy and robustness in noisy or abnormal scenes.
-In accident detection and classification, Yang et al.[9] developed a deep learning model that utilizes multi-vehicle trajectory data for detecting and classifying accidents on freeways, achieving up to 100% detection accuracy.
-For 3D object detection under adverse weather, Qi et al.[10] introduced a novel geometric information constraint network (GIC-Net), enhancing the performance of 3D object detection by incorporating geometric location and line feature constraints, demonstrating superior results under challenging weather conditions.
-In addition to these applications, lane detection has also seen significant advancements, with deep learning-based approaches becoming mainstream due to their strong feature extraction capabilities and generalization.
-Compared to traditional methods, deep learning provides higher robustness, better adapting to complex road environments and effectively handling cases with missing or occluded lane markings.
-LaneNet[11] approaches lane detection as a combination of semantic segmentation and post-processing, first generating a semantic segmentation map of the lanes and then using density clustering to identify each lane instance.
-However, because of the large number of pixels generated during segmentation and the inability of clustering to leverage GPU acceleration, its real-time performance remains limited.
-LaneAF[12], building on the semantic segmentation map, introduced a clustering algorithm based on horizontal and vertical vector fields to accelerate post-processing and used DLA[13] as the backbone network to integrate multi-level semantic and spatial information, enhancing feature representation.
-SCNN[14] leveraged the elongated structure of lane lines by replacing traditional layer-wise convolution with slice convolution, enabling cross-row and cross-column information transfer, which makes it well-suited to line-like lane structures.
-However, its large parameter size prevents real-time performance.
-Overall, current networks, whether for semantic or instance segmentation, mostly rely on pixel-by-pixel classification, though lane detection does not necessarily require dense pixel segmentation.
-To address this, Qin et al.[15] reframed lane detection as a row-based segmentation problem using global features, significantly reducing computational cost and improving detection speed by selecting specific rows and classifying each pixel in the row.
-This method also leverages the large receptive field of global features, enabling it to handle complex scenarios with missing or occluded lanes.
-Apart from segmentation, lane detection also includes detection-based methods, many of which borrow the Anchor-based Methods from object detection[16, 17, 18, 19].
-However, due to the elongated shape of lane lines, traditional bounding boxes are not ideal.
-In lane detection, anchors are often a series of preset starting points and straight lines with different slopes, represented as point sets, with the network predicting the relative distance between the lane line and anchor to optimize line shape and position.
-PointLaneNet[20] first represents lane lines as a series of points, taking the intersections of evenly spaced horizontal lines as coordinates $\{x_{1},x_{2},\ldots,x_{n}\}$ to describe the lane.
-Like other YOLO-based methods[16, 17], PointLaneNet divides the image into multiple grids, predicts lane position and classification at each grid center, and removes duplicate predictions through post-processing.
-However, this grid-based prediction approach has limitations for elongated lanes, reducing detection accuracy.
-LaneATT[21] addresses this issue by noting the high correlation between lane positions and introducing global features to infer lane position, especially in cases of occlusion or missing lane markers.
-LaneATT[21] introduced an anchor-based attention mechanism that significantly improves detection speed and accuracy by utilizing global information.
-CLRNet[22] demonstrated the complementary roles of low- and high-dimensional features in lane detection, proposing Cross Layer Refinement (CLR) to dynamically adjust anchor points through network learning, removing the limitations of fixed anchors to optimize lane detection further.
-FLAMNet[23] proposes an adaptive line anchor framework integrated with a dual-branch network design to improve the accuracy and adaptability of capturing intricate lane geometries.
-By leveraging the Decomposed Self-Attention (DSA) mechanism alongside a Horizontal Feature Fusion Module (HFFM), the model enhances the perceptual range of line anchors in lane detection tasks.
-This approach strengthens robustness against occlusions and incomplete lane markings in challenging environments.
-In addition to these methods, trajectory data has also been explored for lane identification.
-Zhao and Roncoli[24] proposed a constrained spectral clustering method to identify lane information from high-resolution trajectory data, significantly improving results even in the presence of varying lane numbers or lane-changing maneuvers.
-This approach demonstrates the potential of leveraging trajectory data for lane detection, especially in more dynamic environments.
-Compared to other detection and segmentation tasks, research on downstream tasks in adverse weather conditions, such as fog, is relatively limited.
-A conventional solution often applies standard image enhancement techniques, such as classical dehazing methods developed for fog removal and visual quality improvement.
-Nevertheless, enhanced visual clarity doesn’t always correlate with superior detection outcomes.
-Some methods combine image enhancement with detection to reduce adverse weather interference.
-Huang et al.[25] proposed DSNet, incorporating a detection subnet and recovery subnet, where multi-task learning combines dehazing and detection to improve understanding of foggy scenes, resulting in better detection than a standard detection network.
-Additionally, some methods simulate foggy images from existing datasets to optimize downstream tasks.
-For example, Nie et al.[26] used the Monodepth2 model and atmospheric scattering model to convert clear lane images from the CULane dataset into foggy lane images, achieving synthetic data augmentation.
-Testing with SCNN[14] showed that this augmentation strategy improved recognition rates from 76.65% to 86.65% under light fog conditions.
-Another approach employs domain adaptation techniques to address the style gap between foggy scenes and clear-weather images[27, 28].
-Hnewa et al.[29] assumed that there is domain shift between images captured in normal and adverse weather conditions, proposing a multi-scale domain-adaptive YOLO that supports domain adaptation at different levels in the feature extraction stage.
-Li et al.[30] proposed a novel transfer learning-based framework for fog-adaptive object detection, which integrates dual-level adaptation (image-level and object-level) to mitigate domain shifts caused by foggy conditions.
-Notably, the method only requires labeled clear-weather images and unlabeled foggy data for training, eliminating the dependency on annotated foggy datasets.
-Traditional dehazing methods, which are based on statistical or physical principles, often suffer from artifacts such as image darkening or brightening.
-In contrast, end-to-end deep learning methods using Convolutional Neural Networks (CNNs) can capture more complex features but come with higher computational and memory demands.
-These methods are also highly dependent on diverse training datasets.
-However, acquiring real paired images under foggy conditions is challenging, and as a result, many existing datasets rely on synthetic images, which can negatively affect algorithm performance in real-world scenarios.
+Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting.
+Son et al.[3] first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering.
+Hou et al.[4] used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines.
+While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
+With deep learning’s success in image tasks, it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization.
+Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings[5, 6, 7, 8, 9, 10].
+However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation.
+Qin et al.[11] addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
+In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection[12, 13, 14, 15].
+Due to the elongated shape of lane lines, traditional bounding boxes are unsuitable.
+In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position[16, 17, 18, 19, 20, 21, 22].
+These innovations improve accuracy and efficiency in complex or occluded scenarios.
+Research on downstream tasks in adverse weather conditions[23], such as fog, is limited compared to other detection and segmentation tasks.
+Traditional solutions often use image enhancement techniques, like dehazing, to improve visual clarity, but this doesn’t always lead to better detection results.
+Some methods combine image enhancement with detection to reduce weather interference[24], while others simulate foggy images to augment datasets.
+For example, Nie et al.[25] generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
+Domain adaptation techniques also address the style gap between foggy and clear-weather images[26, 27].
+Hnewa et al.[28] proposed a multi-scale domain-adaptive YOLO model, while Li et al.[29] introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
+Traditional dehazing methods suffer from artifacts, whereas deep learning methods using CNNs capture more complex features but require high computational resources and diverse training datasets.
+However, obtaining real paired images under foggy conditions is challenging, and synthetic images often lead to reduced performance in real-world scenarios.
 Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS), supporting key functionalities such as lane departure warnings and autonomous navigation.
@@
 Our method achieves state-of-the-art performance with inference speeds of 192.9 FPS on NVIDIA RTX 3090 and 38.4 FPS on NVIDIA Jetson AGX Orin, confirming its real-time capability and practical applicability in real-world settings.
-2 Related Work
+II Related Work
 In recent years, advances in deep learning have substantially improved lane detection performance.
 Building on these advances, current approaches for lane detection can be categorized into four main paradigms based on their modeling principles: segmentation-based, anchor-based, curve-based, and row-wise-based methods.
-2.1 Segmentation-based Lane Detection
+II-A Segmentation-based Lane Detection
 Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem.
@@
 Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
-2.2 Curve-based Lane Detection
+HW-Transformer[32] presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
+II-B Curve-based Lane Detection
 Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations.
@@
 However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
-2.3 Anchor-based Lane Detection
+PGA-Net[36] introduces an end-to-end framework for lane detection, utilizing cubic polynomial functions and a transformer-based DETR model to simultaneously predict lane shape parameters and incorporate global road context, with the addition of Mean Curvature Loss to enhance the accuracy of curve predictions.
+II-C Anchor-based Lane Detection
 Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings.
@@
 However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
-2.4 Row-wise-based Lane Detection
+II-D Row-wise-based Lane Detection
 Row-wise methods predict lane positions for each image row.
@@
 The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
-3 Data Constrution
+III Data Constrution
 Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models.
@@
 Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[40]
-3.1 FoggyLane
-In the field of intelligent connected vehicle ADAS systems, existing lane detection algorithms have shown good performance under favorable weather conditions.
-However, their performance remains significantly inadequate under adverse weather conditions, such as fog.
-Given that fog is a high-risk scenario for traffic accidents, the limited visibility drastically reduces the driver’s reaction time and decision-making ability, making accidents more likely.
-However, current benchmark datasets for lane detection, including CULane[14], Tusimple[31], and LLAMAS[40], do not include data from foggy scenarios.
-To address this gap, we have developed a lane detection dataset, FoggyLane, specifically designed for foggy conditions.
-Given the randomness of foggy weather, data collection was carried out using a monocular forward-facing camera at different times and locations.
-This includes the morning of October 26, 2023, in Suzhou, Jiangsu Province, China, and two sessions in Urumqi, Xinjiang Uygur Autonomous Region, China: the morning of December 9, 2023, and the night of January 2, 2024.
-In these selected representative traffic scenarios, we sampled one frame every frames, resulting in a total of 1,086 RGB images: 444 images (resolution: $1920\times 1080$), 102 images (resolution: $2560\times 1920$), and 540 images (resolution: $2304\times 1296$).
-To simplify operations and enhance usability during training, we resized all images to a uniform resolution of $1640\times 590$.
-However, since the duration of foggy weather conditions is limited, the number of collected foggy weather images is restricted.
-To improve the model’s performance, we extended our dataset by incorporating additional foggy driving scenes.
-Specifically, we downloaded two foggy driving videos from YouTube[41], extracting road-driving segments at 20-frame intervals, obtaining 44 and 293 frames, respectively.
-This added a total of 337 additional RGB images to the dataset.
-The original resolution of these images was $1920\times 1080$, and they were resized to $1640\times 590$ for consistency.
-As a result, we have constructed FoggyLane, a foggy weather lane detection dataset consisting of 1,423 images.
-FoggyLane includes six types of road scenes, as shown in Fig.
-4, covering Normal, Arrow, Crowd, Curve, Night and Crossroad.
-We split the dataset into training and testing sets at a ratio of 2:1.
-To ensure comprehensive learning of scene characteristics, each type of scene is evenly distributed between the training and testing sets to maintain balance.
-Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
+III-A FoggyLane
+Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents.
+Current benchmark datasets, including CULane[8], Tusimple[30], and LLAMAS[41], lack foggy weather data.
+To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions.
+Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024.
+The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
+Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos[42], resulting in an additional 337 images.
+The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad.
+We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
+Comparing FoggyLane with existing datasets, such as Caltech-Lanes[43] and Tusimple[30], which focus on single traffic scenarios, and LLAMAS[41], which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
+While ApolloScape [44] and BDD100k [45] include multiple traffic scenarios, they lack instance-level annotations.
+Similarly, although CULane [8] and CurveLanes [46] cover a wide range of traffic scenes, they do not include foggy conditions.
+Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced.
+To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple [30] and CULane [8] formats for broader accessibility.
+Figure 4: The different road types of FoggyLane images examples.
+TABLE I: Comparison Between Our Dataset and Existing Lane Detection Datasets
 Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [43] 1.2k None 640 x 480 Tusimple [30] 6.4k None 1280 x 720 LLAMAS [41] 100k None 1276 x 717 Appolloscape [44] 110k None 3384 x 2710 BDD100k [45] 100k None 1280 x 720 CULane [8] 133k None 1640 x 590 CurveLanes [46] 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
-Figure 4: The different road types of FoggyLane images examples.
-Meanwhile, We compared FoggyLane with existing lane detection datasets.
-As shown in Table 1, the Caltech-Lanes and Tusimple datasets only cover single traffic scenarios.
-The LLAMAS [40] dataset focuses on highway scenarios, but its annotations are generated from high-precision maps rather than manually annotated.
-While the ApolloScape [43] and BDD100k [44] datasets contain multiple traffic scenarios, they lack instance-level annotations, and ApolloScape [43] does not provide publicly available test set annotations.
-In comparison, although CULane [14] and CurveLanes [45] include a wide range of traffic scenes with comprehensive annotations, they still lack foggy weather scenarios.
-For lane detection tasks, models trained only on clear weather conditions struggle to handle the challenges posed by fog, such as blurred lane boundaries, faded colors, and reduced image quality.
-To address these limitations, we developed FoggyLane, a comprehensive foggy weather lane detection dataset that includes 1,423 images featuring various challenging scenarios such as foggy roads during the day, curves, crossroads, and night fog.
-The dataset is meticulously annotated using the LabelMe tool, providing instance-level annotations, and it is available in Tusimple [31] and CULane [14] formats for broader accessibility.
 Our dataset has two key advantages:
-  • 1)
+  1. 1.
     Pioneering foggy lane detection dataset: FoggyLane is a dataset specifically designed to address foggy weather scenarios.
@@
 This dataset is crucial for testing the robustness of ADAS under adverse conditions.
-  • 2)
+  2. 2.
     Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[47].
 These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
-Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
-Figure 6: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
-3.2 FoggyCULane and FoggyTusimple
+Figure 5: The FoggyCULane dataset examples.
+III-B FoggyCULane and FoggyTusimple
 Currently, numerous datasets are available for lane detection.
@@
 Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
-The method for generating foggy images is inspired by the model [26], and the complete generation process is shown in Fig.
-5.
-To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions.
-It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one.
-The mathematical modeling equation defined by Koschmieder[47] is represented as:
-$I(x)=J(x)t(x)+A(1-t(x))$ (1)
-where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image.
-$A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
-$t(x)=e^{-\beta d(x)}$ (2)
-where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter.
-$d(x)$ represents the distance between the object and the viewer.
-As this distance increases, the transmittance decreases, resulting in thicker fog.
-Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
-To calculate the corresponding $d(x)$, we use the S2R-DepthNet[48] that has good generalization performance across various applications.
-The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image.
-Then they suppress information irrelevant to depth and use structural information for depth prediction.
-The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information.
-The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image.
-After passing through the DSA Module and STE Module, as shown in Fig.
-5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map.
-The Depth Prediction (DP) is responsible for depth estimation.
-Figure 8: Overview of the proposed network.
-According to the transmittance formula Eq.
-(2), we use the dark channel prior[49] to calculate the average atmospheric light value $A$.
-In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel.
-Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark }}(x)$ can be defined as:
-$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (3)
-where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$.
-In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$.
-After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$.
-The original images and the foggy images generated under different $\beta$ values are displayed in Figs.
-6 and 7.
-4 Methods
-4.1 Overall structure
+The method for generating foggy images is inspired by the model [25], and the specific generation method is provided in the Supplementary Material for Review.
+As shown in Figure 5, it demonstrates the generation case of FoggyCULane.
+IV Methods
+IV-A Overall structure
 To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods.
 As illustrated in the Fig.
-8, the network comprises three main components: a backbone network, a neck network, and a detection head.
+6, the network comprises three main components: a backbone network, a neck network, and a detection head.
 To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network.
@@
 Additionally, to compensate for the loss of edge information in foggy images, we incorporated a low-level edge enhancement module, further improving the network’s ability to detect lane boundaries with precision.
+Figure 6: Overview of the proposed network.
 Figure 7: Backbone Network with Global Feature Fusion.
-4.2 Global Feature Fusion Module
+The Swin Transformer’s window attention mechanism mainly focuses on local features.
+Although its layered structure increases the receptive field, it may still fall short in understanding foggy scenes.
+To address this, we designed a Global Feature Fusion Module (GFFM), replacing the final stage’s window attention with standard multi-head attention to capture global relationships.
+IV-B Global Feature Fusion Module
 In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision.
@@
 Among them, the Swin Transformer[52] combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework.
-By incorporating a sliding window approach similar to convolution, the Swin Transformer limits self-attention calculations within each sliding window, enabling local information interactions while reducing computational load and improving efficiency.
-The Swin Transformer also adopts a multi-scale hierarchical structure similar to that of CNNs, dividing the model into multiple stages.
-In each stage, a Patch Merging layer performs down-sampling to reduce the computational load, followed by the Swin Transformer Block, which applies self-attention within sliding windows.
-The Swin Transformer Block includes both Window Attention and Shifted Window Attention mechanisms; the former computes self-attention within each window, while the latter enables information exchange across windows.
-The Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy.
-Thus, we chose the Swin Transformer as our backbone network.
+As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network.
 However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding.
 Therefore, as illustrated in the Fig.
-9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration.
+7, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration.
 Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs.
 At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
-4.3 Kernel Feature Fusion Module
-Dynamic convolution is a specialized convolution operation that dynamically adjusts kernel parameters based on the input data.
-In conventional convolutions, kernel parameters remain fixed after training and are independent of the input data.
-In contrast, dynamic convolution adapts kernel parameters for each input, enhancing the network’s flexibility and generalization.
-For specific tasks, dynamic convolution often achieves better performance than standard convolution.
-CondInst[57] is an instance segmentation framework based on dynamic convolution, which generates convolution kernel parameters dynamically for each instance based on its features and applies them in the segmentation head to filter the instance’s mask.
-Since each kernel parameter only needs to handle a single instance, this approach reduces the network’s learning burden and convolutional computation load, while also eliminating the ROI operations required in methods like Mask R-CNN, creating a faster, more streamlined instance segmentation framework.
-Other frameworks, such as SOLO[58], SOLOv2[59], and CondLane[38] , have also applied dynamic convolution in instance segmentation, with CondLane specifically designed for lane instance segmentation, achieving robust performance with its dynamic convolution-based lane detection strategy.
-Building on these studies, our lane detection head also incorporates dynamic convolution.
-This detection head includes two components: lane instance detection and lane mask segmentation.
-During instance detection, the network identifies lane instances within the image and generates a set of dynamic convolution kernel parameters for each lane.
-These kernel parameters are applied in the lane mask segmentation head to isolate each lane’s instance segmentation.
-In typical dynamic convolution-based instance segmentation frameworks, the center position features of each instance are used to generate the dynamic convolution kernels.
-However, for lane detection, this is less effective, as lanes are elongated structures, making it challenging for center features to capture information along the full lane length.
-Therefore, following CondLane’s approach, we use the starting position features of each lane to guide the dynamic convolution kernel generation.
-Specifically, we design an instance detection head in the third stage of the neck network’s feature map, comprising two branches: one branch predicts the starting points of lanes, generating a heatmap for these points; the other branch predicts a set of dynamic convolution kernels for each position, using the kernel at the lane’s starting position for subsequent segmentation.
-However, the existing method using lane starting points does not account for the inherent structural and positional relationships among lane instances.
-In a top-down view, lanes are often parallel, similar in shape, and have consistent spacing.
-To fully leverage this structural information, we designed a Kernel Feature Fusion Module (KFFM) to enable the network to automatically learn and predict relationships between lane instances.
-The KFFM is shown in the bottom right of the Fig.
-8.
-4.4 Low-level Edge Enhanced Module
+IV-C Kernel Feature Fusion Module
+Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution.
+CondInst[55] is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency.
+Similarly, CondLane[39] applies dynamic convolution for lane instance segmentation, achieving strong performance.
+Building on this, our lane detection head also uses dynamic convolution, consisting of lane instance detection and mask segmentation.
+Unlike traditional methods, we generate dynamic convolution kernels using the lane’s starting position features rather than center features, which are more suitable for the elongated nature of lanes.
+We design an instance detection head with two branches: one predicts lane starting points and generates a heatmap, and the other generates dynamic convolution kernels for lane segmentation.
+However, existing methods don’t fully utilize the structural relationships between lane instances.
+To address this, we introduce a Kernel Feature Fusion Module (KFFM) that enables the network to learn and predict relationships between lane instances, enhancing detection performance.
+KFFM is shown in the bottom right of Fig.
+6.
+IV-D Low-level Edge Enhanced Module
 Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast.
 This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately.
-Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking.
+Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking[tim1].
 Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig.
-10.
+8.
 It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible.
@@
 Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
+TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
+  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[8] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[11] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[11] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[17] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[17] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[17] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[31] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[31] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[18] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[18] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[18] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[18] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[19] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[19] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
+  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
+* indicates the FPS after acceleration with TensorRT.
 Specifically, we add an auxiliary path in the network to learn the edge features of the input image.
@@
 To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network.
-This branch sequentially applies a $3\times 3$ and a $1\times 1$ convolution.
+This branch sequentially applies a 3 × 3 and a 1 × 1 convolution.
 We then compute the cross-entropy loss between the branch output and the edge labels.
 The detailed structure is illustrated in the bottom-left of the Fig.
-8.
+6.
 Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
-4.5 Loss functions
+Figure 9: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyLane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
+  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[8] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[11] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[11] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[17] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[17] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[17] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[31] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[31] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[18] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[18] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[18] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[18] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[19] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[19] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
+  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
+Meanwhile, in the Dazzle Light scenario, our method is only 0.01 percentage points behind the leading performance.
+Figure 10: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyCULane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+IV-E Loss functions
 Based on the above, our method involves a total of five prediction tasks: lane start point prediction, row-based segmentation prediction, vertical range prediction, offset prediction, and edge prediction.
 The loss functions for each task are detailed below.
-4.5.1 Lane start point prediction
+IV-E1 Lane start point prediction
 In feature maps with a large receptive field and low resolution, the lane line start point is difficult to be accurately represented by a single pixel location.
@@
 The Focal Loss between the predicted heatmap and the ground truth heatmap is defined as:
-$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$ (5)
-4.5.2 Row-based segmentation prediction
+$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$
+IV-E2 Row-based segmentation prediction
 Row-based segmentation predicts a horizontal coordinate $\mathrm{E}\left(\hat{x}_{i}\right)$ for each row.
@@
 $L_{\text{row}}=\frac{1}{N_{v}}\sum_{i\in V}\left|\mathrm{E}\left(\hat{x}_{i}\right)-x_{i}\right|$ (3)
-4.5.3 Vertical range prediction
+IV-E3 Vertical range prediction
 For the label of the lane line vertical range vector prediction, our method uses a $1\times Y$ vector representation.
@@
 $L_{\text{range}}=\sum\left(-y_{gt}^{i}\log(v_{i})-(1-y_{gt}^{i})\log(1-v_{i})\right)$ (4)
-4.5.4 Offset prediction
+IV-E4 Offset prediction
 Firstly, the generation of ground truth labels for offset prediction is required.
@@
 $L_{\text{offset}}=\frac{1}{N_{\Omega}}\sum_{(x,y)\in\Omega}\left|\hat{\delta}_{xy}-\delta_{xy}\right|$ (5)
-4.5.5 Edge prediction
+IV-E5 Edge prediction
 For the task of edge enhancement, the method treats it as an edge segmentation problem, which is essentially a binary classification task.
@@
 $L_{\text{edge}}=\sum_{i}\left(-e_{gt}^{i}\log(p_{i})-(1-e_{gt}^{i})\log(1-p_{i})\right)$ (6)
-4.5.6 Total loss
+IV-E6 Total loss
 To balance the influence of different tasks on the overall training of the network, a weighted fusion of the loss functions for each sub-task is employed.
@@
 where the network achieves optimal performance when the weights are set as $\alpha=1$, $\beta=1$, $\gamma=1$, $\delta=0.4$, and $\varepsilon=0.1$.
-5 Experiments
-5.1 Datasets
+V Experiments
+V-A Datasets
 In the Data Construction chapter, we introduce the FoggyLane dataset, which was constructed from real-world collections for foggy lane detection, along with the FoggyCULane and FoggyTusimple datasets that were synthesized through modeling foggy images.
 Our method is evaluated on these three datasets.
-FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing.
-All images are of resolution $1640\times 590$ pixels.
+FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
 FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach.
 It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing.
-The resolution of all images is also $1640\times 590$ pixels.
+The resolution of all images is also 1640x590 pixels.
 FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method.
 At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases.
-All visual data maintains a consistent resolution of $1280\times 720$ pixels.
-5.2 Evaluation Metric
-Lane detection tasks are typically regarded as a semantic segmentation problem; therefore, the evaluation metrics for lane detection are often similar to those used in segmentation tasks, specifically the confusion matrix: TP (True Positive), FP (False Positive), FN (False Negative), and TN (True Negative).
-In the context of lane detection, TP typically represents the correctly predicted lane line pixels, FP denotes incorrectly predicted lane points, FN refers to lane line points that were not detected by the model but are actually present, and TN indicates correctly predicted background pixels.
-Furthermore, the nature of lane detection tasks introduces some unique characteristics that differentiate their evaluation criteria from those of standard segmentation tasks.
-For FoggyLane and FoggyCULane, we adopt the evaluation metrics of SCNN[14], connecting the labeled or predicted lane points sequentially to form a long line segment with a width of 30 pixels centered around the connecting line.
-We then calculate the Intersection over Union (IoU) between the predicted area and the ground truth area.
-If the IoU of the predicted area and the ground truth area exceeds a certain threshold (typically set at 0.5), the lane detection is considered correct; otherwise, it is deemed incorrect.
-Based on this, we calculate the Precision, Recall, and F1-score, with the following formulas:
+All visual data maintains a consistent resolution of 1280×720 pixels.
+Figure 11: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
+V-B Evaluation Metric
+Lane detection is commonly treated as a semantic segmentation task, evaluated using standard metrics such as True Positive (TP), False Positive (FP), False Negative (FN), and True Negative (TN).
+However, due to its unique structure, lane detection also adopts specialized metrics.
+Following SCNN’s[8] protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas.
+A prediction is considered correct if the Intersection over Union (IoU) with the ground truth exceeds 0.5.
+Based on this, we compute Precision, Recall, and F1-score, with the following formulas:
 $F1=\frac{2\times Precision\times Recall}{Precision+Recall}$ (8)
@@
 Furthermore, if more than 85% of the predicted lane points on a single lane line correspond to the ground truth lane points, the predicted lane is regarded as a true positive.
-5.3 Implementation Details
-We conducted our experiments on the NVIDIA GeForce RTX 3090 platform.
-To achieve better and more stable performance[60], we first pre-trained the SwinGFFM backbone network on the COCO dataset for 500 epochs.
-Subsequently, during the training process for lane detection tasks, we loaded the pre-trained weights of the backbone network and resized the input images to 800 × 320.
-Additionally, we applied various data augmentation techniques to the input images, including rotation, horizontal flipping, brightness adjustment, and random scaling and cropping.
-During training, the batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay strategy.
-For detailed network parameters and training settings, please refer to our open-source code.
-Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
-  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[14] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[15] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[15] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[21] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[21] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[21] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[32] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[32] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[38] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[38] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[38] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[22] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[22] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[22] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[22] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[23] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[23] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
-  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
-* indicates the FPS after acceleration with TensorRT.
-Figure 11: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyLane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
-  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[14] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[15] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[15] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[21] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[21] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[21] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[32] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[32] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[38] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[38] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[38] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[22] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[22] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[22] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[22] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[23] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[23] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
-  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
-Meanwhile, in the Dazzle Light scenario, our method is only 0.01 percentage points behind the leading performance.
-Figure 12: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyCULane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-For objective benchmarking against established lane detection methods including SCNN, RESA, LaneATT, CLRNet, FLAMNet, LSTR, DBNet, UFLD, and CondLane, we replicated these methods on our experimental platform.
-Given the differences in the architecture of each algorithm, the required batch size, number of training epochs, and learning rate vary, and forcing uniform training settings could negatively impact model performance.
-Therefore, we adhered to the training configurations provided in the respective papers.
-Specifically, as the FoggyCULane and FoggyTusimple datasets are extended versions of the CULane and Tusimple datasets, we adopted the original training parameters from CULane and Tusimple when evaluating the algorithms on FoggyCULane and FoggyTusimple.
-For the FoggyLane dataset, we used the training parameters from CULane and extended the training epochs to ensure sufficient training, thereby obtaining reliable test results.
-5.4 Quantitative Results
-5.4.1 Results on FoggyLane
+V-C Implementation Details
+We conducted experiments on the NVIDIA GeForce RTX 3090 platform.
+The SwinGFFM backbone network was pre-trained on the COCO dataset for 500 epochs.
+During lane detection training, we loaded the pre-trained weights and resized input images to 800 × 320.
+Various data augmentation techniques, including rotation, flipping, brightness adjustment, and random scaling, were applied.
+The batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay.
+For full network parameters and training details, please refer to our open-source code.
+To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39], we followed the training settings provided in the respective papers.
+For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
+V-D Quantitative Results
+V-D1 Results on FoggyLane
 We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods.
-As shown in the Tabel 2, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
-From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively.
+As shown in the Tabel II, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
+From the Table II, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively.
 These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%.
@@
 Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
-5.4.2 Results on FoggyCULane
+V-D2 Results on FoggyCULane
 We compared our method with other advanced lane detection methods on the FoggyCULane dataset.
-As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
+As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
 Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
-Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
-5.4.3 Results on FoggyTusimple
-As shown in the Table 4, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods.
+V-D3 Results on FoggyTusimple
+As shown in the Table IV, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods.
 Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%.
 This further demonstrates the effectiveness of our method in foggy conditions.
-Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
+TABLE IV: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
     Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[8] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[11] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[11] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[17] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[17] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[17] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[34] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[31] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[31] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[18] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[18] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[18] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[19] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[19] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[35] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[35] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
-5.5 Ablation Study
-To explore the effectiveness of each part of the model, we conducted ablation studies on both real and synthetic datasets, namely FoggyLane and FoggyCULane.
-We have studied the roles of the GFFM, KFFM, and LEEM modules in the model, with Swin-t as the backbone network.
-As shown in the Tables.
-5 and 6, adding each module boosts the model’s performance on both datasets.
-Among them, GFFM contributes the most to the model’s improvement, with increases of 1.07 on FoggyLane and 0.41 on FoggyCULane.
-This indicates that GFFM can effectively capture both local and global information, enriching the feature representation.
-Building on the foundation of GFFM, the addition of the KFFM module further increases the F1-score by 0.63 on FoggyLane and by 0.26 on FoggyCULane.
-GFFM is capable of automatically learning the associations between lane instance instances, making full use of relevant information.
-On this basis, the introduction of the LEEM module enhances the network’s focus on edge information.
-The F1-score increases by 0.28 on FoggyLane and by 0.14 on FoggyCULane.
-Table 5: Ablation experiments of different components on the FoggyLane dataset.
+V-E Ablation Study
+We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone.
+As shown in Tables V and VI, each module improves the model’s performance.
+The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information.
+Adding the KFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane.
+Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
+TABLE V: Ablation experiments of different components on the FoggyLane dataset.
   Model GFFM KFFM LEEM F1-score Gflops(G)
@@
 Ours 92.34$\uparrow$ 12.01
-Table 6: Ablation experiments of different components on the FoggyCULane dataset.
+TABLE VI: Ablation experiments of different components on the FoggyCULane dataset.
   Model GFFM KFFM LEEM F1-score Gflops(G)
@@
 Ours 78.66$\uparrow$ 12.01
-5.6 Model deployment
+V-F Model deployment
 Figure 12: A edge computing system built with NVIDIA Jetson AGX Orin.
 After experimental validation, it has been demonstrated that our method can run accurately on the vehicle’s edge computing platform with limited computational resources, achieving a performance of 38.4 FPS.
-After considering the deployment of several edge computing tasks[61, 62], we further evaluate the performance of our algorithm in real-world scenarios.
-In addition to experiments on existing datasets, we also deployed and tested the model on the vehicle.
+To further evaluate the performance of our algorithm in real-world scenarios, in addition to experiments on existing datasets, we also deployed and tested the model on the vehicle.
 Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig.
-14.
+12.
 This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit.
 The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
-6 Conclusion
+VI Conclusion
 This study focuses on lane detection in foggy conditions, aiming to address the range of challenges posed by foggy environments, which is crucial for ensuring driving safety in fog and enhancing Advanced Driver Assistance System (ADAS) capabilities.
@@
 2023B1212060029).We would like to release our source code and dataset.
-Compliance with ethics guidelines
-Ronghui Zhang, Yuhang Ma, Tengfei Li, Ziyu Lin, Yueying Wu, Junzhou Chen, Lin Zhang, Jia Hu,Tony Z.
-Qiu and Konghui Guo declare that they have no conflict of interest or financial conflicts to disclose.

### v6

@@
 This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately.
-Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking[tim1].
+Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking.
 Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig.
@@
 TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
-  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[8] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[11] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[11] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[17] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[17] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[17] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[31] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[31] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[18] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[18] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[18] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[18] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[19] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[19] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
+  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[8] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[11] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[11] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[17] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[17] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[17] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[31] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[31] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[18] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[18] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[18] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[18] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[19] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[19] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane[32] IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN[21] IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN[21] IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
   • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
@@
 Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
-Figure 9: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyLane testing set.
+Figure 9: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyLane testing set.
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
 TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
-  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[8] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[11] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[11] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[17] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[17] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[17] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[31] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[31] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[18] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[18] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[18] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[18] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[19] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[19] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.81$\uparrow$)
+  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[8] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[11] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[11] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[17] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[17] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[17] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[31] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[31] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[18] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[18] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[18] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[18] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[19] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[19] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane[32] IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN[21] IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN[21] IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
   • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
-Meanwhile, in the Dazzle Light scenario, our method is only 0.01 percentage points behind the leading performance.
-Figure 10: Visualization results of LaneATT, CondLane, CLRNet, FLAMNet, DBNet, and our method on FoggyCULane testing set.
+Figure 10: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyCULane testing set.
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@
 For full network parameters and training details, please refer to our open-source code.
-To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39], we followed the training settings provided in the respective papers.
+To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39, 32, 21], we followed the training settings provided in the respective papers.
 For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
@@
 As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
-Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
+Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
 V-D3 Results on FoggyTusimple
@@
 TABLE IV: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
-    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[8] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[11] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[11] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[17] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[17] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[17] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[34] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[31] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[31] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[18] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[18] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[18] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[19] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[19] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[35] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[35] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
+    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[8] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[11] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[11] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[17] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[17] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[17] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[34] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[31] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[31] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[18] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[18] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[18] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[19] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[19] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[35] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[35] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane[32] IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN[21] IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN[21] IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
 V-E Ablation Study

### v7

@@
-A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions
+
 Abstract
@@
 Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
-Index Terms:
-ADAS, intelligent vehicle, foggy lane detection dataset, foggy lane detection, global attention, edge enhancement
-I Introduction
+keywords:
+[…]
+1 Introduction
 Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents.
-As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving[1].
+As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Olovsson et al., 2022).
 Under good weather and clear lane markings, current lane detection algorithms perform well.
@@
 This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
-Figure 1: Lane detection in foggy scenarios for ADAS.[2]
+Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2).
+This is partly due to the limited development of connected vehicle infrastructure and data platforms, which constrains the large-scale deployment of higher-level autonomous functions (Lim et al., 2021).
+In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments.
+Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions.
+Research related to vehicle trajectory analysis (Berghaus et al., 2024), congestion-aware traffic prediction (Zhang et al., 2025b), and 3D scene understanding (Qu et al., 2025) has become essential for improving system performance and safety in real-world driving environments, providing valuable insights for the further enhancement of lane detection capabilities in these challenging scenarios.
+Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels, 2025)
 Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting.
-Son et al.[3] first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering.
-Hou et al.[4] used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines.
+Son et al.(Son et al., 2015) first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering.
+Hou et al.(Hou et al., 2016) used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines.
 While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
 With deep learning’s success in image tasks, it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization.
-Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings[5, 6, 7, 8, 9, 10].
+Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018).
 However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation.
-Qin et al.[11] addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
-In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection[12, 13, 14, 15].
+Qin et al.(Qin et al., 2020) addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
+In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection(Redmon, 2016; Farhadi and Redmon, 2018).
 Due to the elongated shape of lane lines, traditional bounding boxes are unsuitable.
-In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position[16, 17, 18, 19, 20, 21, 22].
+In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position(Chen et al., 2019; Zheng et al., 2022; Ran et al., 2023; Wang et al., 2025).
 These innovations improve accuracy and efficiency in complex or occluded scenarios.
-Research on downstream tasks in adverse weather conditions[23], such as fog, is limited compared to other detection and segmentation tasks.
+Research on downstream tasks in adverse weather conditions, such as fog, is limited compared to other detection and segmentation tasks.
 Traditional solutions often use image enhancement techniques, like dehazing, to improve visual clarity, but this doesn’t always lead to better detection results.
-Some methods combine image enhancement with detection to reduce weather interference[24], while others simulate foggy images to augment datasets.
-For example, Nie et al.[25] generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
-Domain adaptation techniques also address the style gap between foggy and clear-weather images[26, 27].
-Hnewa et al.[28] proposed a multi-scale domain-adaptive YOLO model, while Li et al.[29] introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
+Some methods combine image enhancement with detection to reduce weather interference(Huang et al., 2020), while others simulate foggy images to augment datasets.
+For example, Nie et al.(Nie et al., 2022) generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
+Domain adaptation techniques also address the style gap between foggy and clear-weather images(Chen et al., 2018; Zhang et al., 2021).
+Hnewa et al.(Hnewa and Radha, 2021) proposed a multi-scale domain-adaptive YOLO model, while Li et al.(Li et al., 2023) introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
 Traditional dehazing methods suffer from artifacts, whereas deep learning methods using CNNs capture more complex features but require high computational resources and diverse training datasets.
@@
 This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets.
-In addition, we generate foggy versions of two widely-used lane detection datasets, CULane [8] and Tusimple [30], creating FoggyCULane and FoggyTusimple datasets.
+In addition, we generate foggy versions of two widely-used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), creating FoggyCULane and FoggyTusimple datasets.
 These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@
 Our method achieves state-of-the-art performance with inference speeds of 192.9 FPS on NVIDIA RTX 3090 and 38.4 FPS on NVIDIA Jetson AGX Orin, confirming its real-time capability and practical applicability in real-world settings.
-II Related Work
+2 Related Work
 In recent years, advances in deep learning have substantially improved lane detection performance.
 Building on these advances, current approaches for lane detection can be categorized into four main paradigms based on their modeling principles: segmentation-based, anchor-based, curve-based, and row-wise-based methods.
-II-A Segmentation-based Lane Detection
+2.1 Segmentation-based Lane Detection
 Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem.
-LaneNet[5] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines.
-SCNN[8] extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition.
-RESA[31] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information.
+LaneNet(Neven et al., 2018) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines.
+SCNN(Pan et al., 2018) extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition.
+RESA(Zheng et al., 2021) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information.
 Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
-HW-Transformer[32] presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
-II-B Curve-based Lane Detection
+HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
+2.2 Curve-based Lane Detection
 Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations.
-PolyLaneNet[33] introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane.
-LSTR[34] can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information.
-DBNet[35] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation.
+PolyLaneNet(Tabelini et al., 2021b) introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane.
+LSTR(Liu et al., 2021b) can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information.
+DBNet(Dai et al., 2024) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation.
 However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
-PGA-Net[36] introduces an end-to-end framework for lane detection, utilizing cubic polynomial functions and a transformer-based DETR model to simultaneously predict lane shape parameters and incorporate global road context, with the addition of Mean Curvature Loss to enhance the accuracy of curve predictions.
-II-C Anchor-based Lane Detection
+2.3 Anchor-based Lane Detection
 Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings.
-Line-CNN[37] introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors.
-LaneATT[17] effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module.
-CLRNet[18] employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features.
+Line-CNN(Li et al., 2019) introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors.
+LaneATT(Tabelini et al., 2021a) effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module.
+CLRNet(Zheng et al., 2022) employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features.
 This hierarchical fusion of semantic information significantly improves detection precision.
-FLAMNet[19] employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms.
+FLAMNet(Ran et al., 2023) employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms.
 This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures.
 However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
-II-D Row-wise-based Lane Detection
+2.4 Row-wise-based Lane Detection
 Row-wise methods predict lane positions for each image row.
-E2E-LMD[38] first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions.
-UFLD[11] utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes.
-CondLaneNet[39] utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts.
+E2E-LMD(Yoo et al., 2020) first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions.
+UFLD(Qin et al., 2020) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes.
+CondLaneNet(Liu et al., 2021a) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts.
 The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
-III Data Constrution
+3 Data Constrution
 Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models.
@@
 To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane.
-In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [8] and Tusimple [30] datasets to test the generalizability of our methods.
-Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.[40]
-Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[40]
-III-A FoggyLane
+In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017) datasets to test the generalizability of our methods.
+Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(GoogleMaps, Accessed: 2024)
+Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(GoogleMaps, Accessed: 2024)
+3.1 FoggyLane
 Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents.
-Current benchmark datasets, including CULane[8], Tusimple[30], and LLAMAS[41], lack foggy weather data.
+Current benchmark datasets, including CULane(Pan et al., 2018), Tusimple(Tusimple, 2017), and LLAMAS(Behrendt and Soussan, 2019), lack foggy weather data.
 To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions.
@@
 The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
-Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos[42], resulting in an additional 337 images.
+Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube, Accessed: 2024), resulting in an additional 337 images.
 The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad.
 We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
-Comparing FoggyLane with existing datasets, such as Caltech-Lanes[43] and Tusimple[30], which focus on single traffic scenarios, and LLAMAS[41], which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
-While ApolloScape [44] and BDD100k [45] include multiple traffic scenarios, they lack instance-level annotations.
-Similarly, although CULane [8] and CurveLanes [46] cover a wide range of traffic scenes, they do not include foggy conditions.
+Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
+While ApolloScape (Huang et al., 2018) and BDD100k (Yu et al., 2020) include multiple traffic scenarios, they lack instance-level annotations.
+Similarly, although CULane (Pan et al., 2018) and CurveLanes (Xu et al., 2020) cover a wide range of traffic scenes, they do not include foggy conditions.
 Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced.
-To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple [30] and CULane [8] formats for broader accessibility.
+To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple (Tusimple, 2017) and CULane (Pan et al., 2018) formats for broader accessibility.
 Figure 4: The different road types of FoggyLane images examples.
-TABLE I: Comparison Between Our Dataset and Existing Lane Detection Datasets
-Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [43] 1.2k None 640 x 480 Tusimple [30] 6.4k None 1280 x 720 LLAMAS [41] 100k None 1276 x 717 Appolloscape [44] 110k None 3384 x 2710 BDD100k [45] 100k None 1280 x 720 CULane [8] 133k None 1640 x 590 CurveLanes [46] 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
+Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
+Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
 Our dataset has two key advantages:
+  1. []
   2. 1.
@@
   3. 2.
-    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[47].
+    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.(Zhang et al., 2024a).
 These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
+Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
 Figure 6: The FoggyCULane dataset examples.
-III-B FoggyCULane and FoggyTusimple
+3.2 FoggyCULane and FoggyTusimple
 Currently, numerous datasets are available for lane detection.
-Among them, we select two widely adopted and representative benchmarks: CULane[8] and TuSimple[30].
-The CULane dataset [8] is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing.
-In contrast, the TuSimple dataset [30] primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
+Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017).
+The CULane dataset (Pan et al., 2018) is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing.
+In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
 Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
-The method for generating foggy images is inspired by the model [25], and the specific generation method is provided in the Supplementary Material for Review.
-As shown in Figure 5, it demonstrates the generation case of FoggyCULane.
-IV Methods
-IV-A Overall structure
+The method for generating foggy images is inspired by the model (Nie et al., 2022), and the complete generation process is shown in Fig.
+5.
+To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions.
+It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one.
+The mathematical modeling equation defined by Koschmieder (Harald, 1924) is represented as:
+$I(x)=J(x)t(x)+A(1-t(x))$ (1)
+where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image.
+$A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
+$t(x)=e^{-\beta d(x)}$ (2)
+where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter.
+$d(x)$ represents the distance between the object and the viewer.
+As this distance increases, the transmittance decreases, resulting in thicker fog.
+Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
+To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications.
+The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image.
+Then they suppress information irrelevant to depth and use structural information for depth prediction.
+The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information.
+The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image.
+After passing through the DSA Module and STE Module, as shown in Fig.
+5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map.
+The Depth Prediction (DP) is responsible for depth estimation.
+According to the transmittance formula Eq.
+(2), we use the dark channel prior(He et al., 2011) to calculate the average atmospheric light value $A$.
+In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel.
+Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
+$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (3)
+where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$.
+In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$.
+After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$.
+The original images and the foggy images generated under different $\beta$ values are displayed in Figs.
+6 and 7.
+4 Methods
+4.1 Overall structure
 To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods.
 As illustrated in the Fig.
-6, the network comprises three main components: a backbone network, a neck network, and a detection head.
+8, the network comprises three main components: a backbone network, a neck network, and a detection head.
 To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network.
-In the neck network, we utilize HRNetV2[48] to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
+In the neck network, we utilize HRNetV2(Wang et al., 2021) to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
 For the detection head, we employ the concept of dynamic convolution, incorporating both lane instance detection and lane mask segmentation.
@@
 Figure 9: Backbone Network with Global Feature Fusion.
-The Swin Transformer’s window attention mechanism mainly focuses on local features.
-Although its layered structure increases the receptive field, it may still fall short in understanding foggy scenes.
-To address this, we designed a Global Feature Fusion Module (GFFM), replacing the final stage’s window attention with standard multi-head attention to capture global relationships.
-IV-B Global Feature Fusion Module
+4.2 Global Feature Fusion Module
 In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision.
 With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information.
-As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[49, 50, 51].
+As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025a; Chen et al., 2023; Zhang et al., 2024b).
 However, applying Transformers to the visual domain still faces some challenges.
@@
 Additionally, calculating global attention at high resolutions can be computationally intensive.
-To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers[52, 53, 54].
-Among them, the Swin Transformer[52] combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework.
+To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers(Liu et al., 2021c, 2022; Dong et al., 2022).
+Among them, the Swin Transformer(Liu et al., 2021c) combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework.
 As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network.
@@
 Therefore, as illustrated in the Fig.
-7, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration.
+9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration.
 Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs.
 At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
-IV-C Kernel Feature Fusion Module
+4.3 Kernel Feature Fusion Module
 Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution.
-CondInst[55] is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency.
-Similarly, CondLane[39] applies dynamic convolution for lane instance segmentation, achieving strong performance.
+CondInst(Tian et al., 2023) is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency.
+Similarly, CondLane(Liu et al., 2021a) applies dynamic convolution for lane instance segmentation, achieving strong performance.
 Building on this, our lane detection head also uses dynamic convolution, consisting of lane instance detection and mask segmentation.
@@
 KFFM is shown in the bottom right of Fig.
-6.
-IV-D Low-level Edge Enhanced Module
+8.
+4.4 Low-level Edge Enhanced Module
 Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast.
@@
 Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig.
-8.
+10.
 It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible.
@@
 Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
-TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
-  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[8] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[11] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[11] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[17] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[17] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[17] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[31] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[31] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[18] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[18] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[18] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[18] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[19] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[19] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane[32] IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN[21] IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN[21] IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
+Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
+  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN(Pan et al., 2018) AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
   • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
@@
 The detailed structure is illustrated in the bottom-left of the Fig.
-6.
+8.
 Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
-Figure 9: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyLane testing set.
+Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set.
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
-  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[8] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[11] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[11] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[17] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[17] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[17] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[31] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[31] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[18] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[18] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[18] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[18] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[19] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[19] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane[32] IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN[21] IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN[21] IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
-  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
-Figure 10: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyCULane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-IV-E Loss functions
+4.5 Loss functions
 Based on the above, our method involves a total of five prediction tasks: lane start point prediction, row-based segmentation prediction, vertical range prediction, offset prediction, and edge prediction.
 The loss functions for each task are detailed below.
-IV-E1 Lane start point prediction
+4.5.1 Lane start point prediction
 In feature maps with a large receptive field and low resolution, the lane line start point is difficult to be accurately represented by a single pixel location.
@@
 The Focal Loss between the predicted heatmap and the ground truth heatmap is defined as:
-$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$
-IV-E2 Row-based segmentation prediction
+$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$ (5)
+4.5.2 Row-based segmentation prediction
 Row-based segmentation predicts a horizontal coordinate $\mathrm{E}\left(\hat{x}_{i}\right)$ for each row.
@@
 $L_{\text{row}}=\frac{1}{N_{v}}\sum_{i\in V}\left|\mathrm{E}\left(\hat{x}_{i}\right)-x_{i}\right|$ (6)
-IV-E3 Vertical range prediction
+4.5.3 Vertical range prediction
 For the label of the lane line vertical range vector prediction, our method uses a $1\times Y$ vector representation.
@@
 $L_{\text{range}}=\sum\left(-y_{gt}^{i}\log(v_{i})-(1-y_{gt}^{i})\log(1-v_{i})\right)$ (7)
-IV-E4 Offset prediction
+4.5.4 Offset prediction
 Firstly, the generation of ground truth labels for offset prediction is required.
@@
 $L_{\text{offset}}=\frac{1}{N_{\Omega}}\sum_{(x,y)\in\Omega}\left|\hat{\delta}_{xy}-\delta_{xy}\right|$ (8)
-IV-E5 Edge prediction
+4.5.5 Edge prediction
 For the task of edge enhancement, the method treats it as an edge segmentation problem, which is essentially a binary classification task.
@@
 $L_{\text{edge}}=\sum_{i}\left(-e_{gt}^{i}\log(p_{i})-(1-e_{gt}^{i})\log(1-p_{i})\right)$ (9)
-IV-E6 Total loss
+4.5.6 Total loss
 To balance the influence of different tasks on the overall training of the network, a weighted fusion of the loss functions for each sub-task is employed.
@@
 where the network achieves optimal performance when the weights are set as $\alpha=1$, $\beta=1$, $\gamma=1$, $\delta=0.4$, and $\varepsilon=0.1$.
-V Experiments
-V-A Datasets
+Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
+  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN(Pan et al., 2018) AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
+  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
+Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+5 Experiments
+5.1 Datasets
 In the Data Construction chapter, we introduce the FoggyLane dataset, which was constructed from real-world collections for foggy lane detection, along with the FoggyCULane and FoggyTusimple datasets that were synthesized through modeling foggy images.
@@
 Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
-V-B Evaluation Metric
+5.2 Evaluation Metric
 Lane detection is commonly treated as a semantic segmentation task, evaluated using standard metrics such as True Positive (TP), False Positive (FP), False Negative (FN), and True Negative (TN).
 However, due to its unique structure, lane detection also adopts specialized metrics.
-Following SCNN’s[8] protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas.
+Following SCNN’s(Pan et al., 2018) protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas.
 A prediction is considered correct if the Intersection over Union (IoU) with the ground truth exceeds 0.5.
@@
 Furthermore, if more than 85% of the predicted lane points on a single lane line correspond to the ground truth lane points, the predicted lane is regarded as a true positive.
-V-C Implementation Details
+Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
+    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN(Pan et al., 2018) AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
+5.3 Implementation Details
 We conducted experiments on the NVIDIA GeForce RTX 3090 platform.
@@
 For full network parameters and training details, please refer to our open-source code.
-To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39, 32, 21], we followed the training settings provided in the respective papers.
+To benchmark against existing methods(Pan et al., 2018; Zheng et al., 2021; Tabelini et al., 2021a; Zheng et al., 2022; Ran et al., 2023; Liu et al., 2021b; Dai et al., 2024; Qin et al., 2020; Liu et al., 2021a; Wang et al., 2025), we followed the training settings provided in the respective papers.
 For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
-V-D Quantitative Results
-V-D1 Results on FoggyLane
+5.4 Quantitative Results
+5.4.1 Results on FoggyLane
 We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods.
-As shown in the Tabel II, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
-From the Table II, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively.
+As shown in the Tabel 2, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
+From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively.
 These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%.
@@
 Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
-V-D2 Results on FoggyCULane
+5.4.2 Results on FoggyCULane
 We compared our method with other advanced lane detection methods on the FoggyCULane dataset.
-As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
+As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
 Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
-V-D3 Results on FoggyTusimple
-As shown in the Table IV, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods.
+5.4.3 Results on FoggyTusimple
+As shown in the Table 4, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods.
 Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%.
 This further demonstrates the effectiveness of our method in foggy conditions.
-TABLE IV: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
-    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[8] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[11] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[11] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[17] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[17] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[17] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[34] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[31] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[31] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[18] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[18] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[18] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[19] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[19] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[35] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[35] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane[32] IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN[21] IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN[21] IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
-V-E Ablation Study
+5.5 Ablation Study
 We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone.
-As shown in Tables V and VI, each module improves the model’s performance.
+As shown in Tables 5 and 6, each module improves the model’s performance.
 The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information.
@@
 Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
-TABLE V: Ablation experiments of different components on the FoggyLane dataset.
+Table 5: Ablation experiments of different components on the FoggyLane dataset.
   Model GFFM KFFM LEEM F1-score Gflops(G)
@@
 Ours 92.34$\uparrow$ 12.01
-TABLE VI: Ablation experiments of different components on the FoggyCULane dataset.
+Table 6: Ablation experiments of different components on the FoggyCULane dataset.
   Model GFFM KFFM LEEM F1-score Gflops(G)
@@
 Ours 78.66$\uparrow$ 12.01
-V-F Model deployment
+5.6 Model deployment
 Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin.
@@
 Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig.
-12.
+14.
 This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit.
 The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
-VI Conclusion
+6 Conclusion
 This study focuses on lane detection in foggy conditions, aiming to address the range of challenges posed by foggy environments, which is crucial for ensuring driving safety in fog and enhancing Advanced Driver Assistance System (ADAS) capabilities.

### v8

@@
 
-Abstract
-Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS).
-Existing lane detection algorithms generally perform well under favorable weather conditions.
-However, their performance degrades significantly in adverse conditions, such as fog, which increases the risk of traffic accidents.
-This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments.
-To address this, we introduce the FoggyLane dataset, captured in real-world foggy scenarios, and synthesize two additional datasets, FoggyCULane and FoggyTusimple, from existing popular lane detection datasets.
-Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions.
-Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, with F1-scores of 95.04 on FoggyLane, 79.85 on FoggyCULane, and 96.95 on FoggyTusimple.
-Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
-keywords:
+[…]
 […]
@@
 Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents.
-As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Olovsson et al., 2022).
+As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Ma and Xue, 2024).
 Under good weather and clear lane markings, current lane detection algorithms perform well.
-However, in complex road environments and poor weather, such as fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
+However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
 In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences.
@@
 Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2).
-This is partly due to the limited development of connected vehicle infrastructure and data platforms, which constrains the large-scale deployment of higher-level autonomous functions (Lim et al., 2021).
+This is partly due to the limited development of connected vehicle infrastructure and data platforms, which constrains the large-scale deployment of higher-level autonomous functions (Qu et al., 2025).
 In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments.
 Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions.
-Research related to vehicle trajectory analysis (Berghaus et al., 2024), congestion-aware traffic prediction (Zhang et al., 2025b), and 3D scene understanding (Qu et al., 2025) has become essential for improving system performance and safety in real-world driving environments, providing valuable insights for the further enhancement of lane detection capabilities in these challenging scenarios.
+Research related to vehicle trajectory analysis (Berghaus et al., 2024) and 3D scene understanding (Lim et al., 2021) has become essential for improving system performance and safety in real-world driving environments, providing valuable insights for the further enhancement of lane detection capabilities in these challenging scenarios.
 Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels, 2025)
@@
 While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
-With deep learning’s success in image tasks, it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization.
+With deep learning’s success in many traffic tasks(Guo et al., 2023; Yang et al., 2023a; Yuan et al., 2023; Zaman et al., 2025), it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization.
 Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018).
@@
 We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
+Figure 4: The different road types of FoggyLane images examples.
+Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
+Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
 Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
@@
 To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple (Tusimple, 2017) and CULane (Pan et al., 2018) formats for broader accessibility.
-Figure 4: The different road types of FoggyLane images examples.
-Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
-Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
 Our dataset has two key advantages:
-  1. []
   1. 1.
@@
 These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
-Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
-Figure 6: The FoggyCULane dataset examples.
 3.2 FoggyCULane and FoggyTusimple
@@
 Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
+Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
 The method for generating foggy images is inspired by the model (Nie et al., 2022), and the complete generation process is shown in Fig.
@@
 As this distance increases, the transmittance decreases, resulting in thicker fog.
-Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
+Figure 6: The FoggyCULane dataset examples.
 To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications.
@@
 The Depth Prediction (DP) is responsible for depth estimation.
+Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
 According to the transmittance formula Eq.
@@
 In the neck network, we utilize HRNetV2(Wang et al., 2021) to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
+Figure 8: Overview of the proposed network.
 For the detection head, we employ the concept of dynamic convolution, incorporating both lane instance detection and lane mask segmentation.
@@
 Additionally, to compensate for the loss of edge information in foggy images, we incorporated a low-level edge enhancement module, further improving the network’s ability to detect lane boundaries with precision.
-Figure 8: Overview of the proposed network.
-Figure 9: Backbone Network with Global Feature Fusion.
 4.2 Global Feature Fusion Module
@@
 With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information.
-As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025a; Chen et al., 2023; Zhang et al., 2024b).
+As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025b; Chen et al., 2023; Zhang et al., 2024b; Yang et al., 2023b; Zhang et al., 2025a).
 However, applying Transformers to the visual domain still faces some challenges.
@@
 Additionally, calculating global attention at high resolutions can be computationally intensive.
+Figure 9: Backbone Network with Global Feature Fusion.
 To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers(Liu et al., 2021c, 2022; Dong et al., 2022).
@@
 Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
-Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
-  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN(Pan et al., 2018) AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
-  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
-* indicates the FPS after acceleration with TensorRT.
 Specifically, we add an auxiliary path in the network to learn the edge features of the input image.
@@
 Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
-Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
 4.5 Loss functions
@@
 where the network achieves optimal performance when the weights are set as $\alpha=1$, $\beta=1$, $\gamma=1$, $\delta=0.4$, and $\varepsilon=0.1$.
-Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
-  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN(Pan et al., 2018) AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
-  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow, Curve and Night conditions, as well as in the overall (Total) performance.
-Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
 5 Experiments
@@
 All visual data maintains a consistent resolution of 1280×720 pixels.
-Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
+Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
+  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN(Pan et al., 2018) AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.6 147.6 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.6 70.0 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.0 171.6 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.0 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(4.63$\uparrow$) 77.32(5.17$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.32$\uparrow$) 20.3 38.2 / 192.9*
+  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
+* indicates the FPS after acceleration with TensorRT.
+Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
 5.2 Evaluation Metric
@@
 Furthermore, if more than 85% of the predicted lane points on a single lane line correspond to the ground truth lane points, the predicted lane is regarded as a true positive.
-Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
-    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN(Pan et al., 2018) AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
 5.3 Implementation Details
@@
 To benchmark against existing methods(Pan et al., 2018; Zheng et al., 2021; Tabelini et al., 2021a; Zheng et al., 2022; Ran et al., 2023; Liu et al., 2021b; Dai et al., 2024; Qin et al., 2020; Liu et al., 2021a; Wang et al., 2025), we followed the training settings provided in the respective papers.
-For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
+For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
+Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
+  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN(Pan et al., 2018) AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.07$\uparrow$) 78.63(0.87$\uparrow$) 73.89 82.74(1.85$\uparrow$) 53.75(0.50$\uparrow$) 89.82(0.37$\uparrow$) 74.09 1146 74.97(0.76$\uparrow$) 79.85(0.72$\uparrow$)
+  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow and Night conditions, as well as in the overall (Total) performance.
+Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
 5.4 Quantitative Results
@@
 From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively.
-These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%.
-The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach.
+These results surpass the best-performing alternative methods by margins of 2.95%, 4.63%, 5.17%, and 3.41%.
+The mean F1 (mF1) score demonstrates a 3.32% improvement over the strongest competing approach.
 Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
@@
 We compared our method with other advanced lane detection methods on the FoggyCULane dataset.
-As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve.
+As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow and Curve.
 Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
@@
 This further demonstrates the effectiveness of our method in foggy conditions.
+Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
+    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN(Pan et al., 2018) AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.89$\downarrow$) 3.83
+Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
 5.5 Ablation Study
@@
 Ours 92.34$\uparrow$ 12.01
+5.6 Model deployment
+To further evaluate the performance of our algorithm in real-world scenarios, in addition to experiments on existing datasets, we also deployed and tested the model on the vehicle.
+Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig.
+14.
+This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit.
+The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
 Table 6: Ablation experiments of different components on the FoggyCULane dataset.
@@
 Ours 78.66$\uparrow$ 12.01
-5.6 Model deployment
 Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin.
 After experimental validation, it has been demonstrated that our method can run accurately on the vehicle’s edge computing platform with limited computational resources, achieving a performance of 38.4 FPS.
-To further evaluate the performance of our algorithm in real-world scenarios, in addition to experiments on existing datasets, we also deployed and tested the model on the vehicle.
-Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig.
-14.
-This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit.
-The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
 6 Conclusion
@@
 Furthermore, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, further validating its practical applicability in real-world scenarios.
-Acknowledgments
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgement
 This project is jointly supported by National Natural Science Foundation of China (Nos.
@@
 2023B1212060029).We would like to release our source code and dataset.
+Author biography
+{biography}
+[zrh.jpg]Ronghui Zhang received a B.Sc.
+(Eng.) from the Department of Automation Science and Electrical Engineering, Hebei University, Baoding, China, in 2003, an M.S.
+degree in Vehicle Application Engineering from Jilin University, Changchun, China, in 2006, and a Ph.D.
+(Eng.) in Mechanical & Electrical Engineering from Changchun Institute of Optics, Fine Mechanics and Physics, the Chinese Academy of Sciences, Changchun, China, in 2009.
+After finishing his post-doctoral research work at INRIA, Paris, France, in February 2011, he is currently an Associate Professor with the Research Center of Intelligent Transportation Systems, School of intelligent systems engineering, Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China.
+His current research interests include computer vision, intelligent control and ITS.
+{biography}
+[mayh.jpg]Yuhang Ma received his B.S.
+degree in Automation at Huazhong University of Science and Technology in 2020.
+He is currently pursuing the master’s degree in Electronic and Information Engineering at Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China.
+His current research interests include autonomous driving and computer vision.
+{biography}
+[ltf.jpg]Tengfei Li received his B.S.
+degree in Robot Engineering from Harbin Institute of Technology at Weihai in 2024.
+He is currently pursuing the M.S.
+degree with the School of Intelligent Systems Engineering, Sun Yat-sen University.
+His research interests include computer vision and autonomous driving.
+{biography}[linziyu.pdf]Ziyu Lin is currently pursuing a B.Sc.
+degree in Traffic Engineering with the School of intelligent systems engineering, Sun Yat-sen University, Guangzhou, China.
+Her research interests include computer vision, deep learning, and autonomous driving technology, with a particular focus on exploring how these technologies can enhance the efficiency and safety of autonomous driving.
+{biography}
+[wyy.jpg]Yueying Wu received her B.S.
+degree in Internet of Things Engineering from Central South University, Changsha, China, in 2020, and an M.S.
+degree in Transportation Engineering from Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China, in 2023.
+Her current research interests include autonomous driving and vehicle-road collaboration, deep learning and computer vision.
+{biography}[cjz.png]Junzhou Chen received his Ph.D.
+in Computer Science and Engineering from the Chinese University of Hong Kong in 2008, following his M.Eng degree in Software Engineering and B.S.
+in Computer Science and Applications from Sichuan University in 2005 and 2002, respectively.
+Between March 2009 and February 2019, he served as a Lecturer and later as an Associate Professor at the School of Information Science and Technology at Southwest Jiaotong University.
+He is currently an associate professor at the School of Intelligent Systems Engineering at Sun Yat-sen University.
+His research interests include computer vision, machine learning, intelligent transportation systems, mobile computing and medical image processing.
+{biography}
+[…]

### v9

@@
-
+A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions
+Abstract
+Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS).
+Existing lane detection algorithms generally perform well under favorable weather conditions.
+However, their performance degrades significantly in adverse conditions, such as fog, which increases the risk of traffic accidents.
+This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments.
+To address this, we introduce the FoggyLane dataset, captured in real-world foggy scenarios, and synthesize two additional datasets, FoggyCULane and FoggyTusimple, from existing popular lane detection datasets.
+Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions.
+Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, with F1-scores of 95.04 on FoggyLane, 79.85 on FoggyCULane, and 96.95 on FoggyTusimple.
+Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
+keywords:
+ADAS, Intelligent vehicle, Foggy lane detection dataset, Foggy lane detection, Global attention, Edge enhancement
+journal: Expert Systems With Applications
 […]
+[1]organization=Guangdong Key Laboratory of Intelligent Transportation System, School of Intelligent Systems Engineering, Sun Yat-sen University, city=Guangzhou, postcode=510275, country=China
 […]
+[2]organization=College of Automotive Studies, Tongji University, city=Shanghai, postcode=201804, country=China
+[…]
+[3]organization=College of Transportation Engineering, Tongji University, city=Shanghai, postcode=201804, country=China
+[…]
+[4]organization=Department of Civil and Environmental Engineering, University of Alberta, city=Edmonton, state=Alberta, country=Canada
+[…]
+[5]organization=State Key Laboratory of Automotive Chassis Integration and Bionics, Jilin University, city=Changchun, postcode=130025, country=China
 1 Introduction
 Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents.
-As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Ma and Xue, 2024).
+As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Chen et al., 2026).
 Under good weather and clear lane markings, current lane detection algorithms perform well.
-However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
+However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging(Sang and Norris, 2025).
 In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences.
@@
 This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
+Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels, 2025)
 Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2).
-This is partly due to the limited development of connected vehicle infrastructure and data platforms, which constrains the large-scale deployment of higher-level autonomous functions (Qu et al., 2025).
 In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments.
 Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions.
-Research related to vehicle trajectory analysis (Berghaus et al., 2024) and 3D scene understanding (Lim et al., 2021) has become essential for improving system performance and safety in real-world driving environments, providing valuable insights for the further enhancement of lane detection capabilities in these challenging scenarios.
-Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels, 2025)
-Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting.
+For example, Zhang and Zhu (2019) proposed an autonomous path tracking control method for intelligent electric vehicles based on lane detection and an optimal preview strategy, demonstrating that accurate lane recognition and trajectory tracking are fundamental to ensuring driving stability and safety at lower automation levels (Zhang and Zhu, 2019).
+Similarly, recent research highlights that coordinated lane-changing strategies in mixed traffic, such as the decentralized multi-vehicle decision-making framework developed by Alshakhs et al., are essential for improving both traffic efficiency and safety in environments where human-driven vehicles and connected automated vehicles coexist (Alshakhs et al., 2025).
+Therefore, this research not only focuses on end-to-end autonomous driving technologies but also aims to improve lane detection performance in complex environments, providing stronger safety support and driving assistance for L0, L1, and L2 vehicles in adverse conditions.
+Earlier lane detection algorithms mainly relied on traditional computer vision methods.
 Son et al.(Son et al., 2015) first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering.
@@
 While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
-With deep learning’s success in many traffic tasks(Guo et al., 2023; Yang et al., 2023a; Yuan et al., 2023; Zaman et al., 2025), it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization.
+With deep learning’s success in many traffic tasks(Zhang et al., 2024b; Shunmuga Perumal et al., 2023; Preeti and Rana, 2024; Awan et al., 2025), it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization.
 Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018).
@@
 However, obtaining real paired images under foggy conditions is challenging, and synthetic images often lead to reduced performance in real-world scenarios.
-Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS), supporting key functionalities such as lane departure warnings and autonomous navigation.
+Lane detection enables key functionalities such as lane departure warnings and autonomous navigation in ADAS.
 While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents.
@@
 This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets.
-In addition, we generate foggy versions of two widely-used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), creating FoggyCULane and FoggyTusimple datasets.
+Additionally, we create foggy versions of two commonly used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), resulting in the FoggyCULane and FoggyTusimple datasets.
 These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@
 In recent years, advances in deep learning have substantially improved lane detection performance.
-Building on these advances, current approaches for lane detection can be categorized into four main paradigms based on their modeling principles: segmentation-based, anchor-based, curve-based, and row-wise-based methods.
+Building on these advances, current lane detection methods are divided into four main paradigms: segmentation-based, anchor-based, curve-based, and row-wise-based approaches, according to their modeling strategies.
 2.1 Segmentation-based Lane Detection
@@
 LaneNet(Neven et al., 2018) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines.
-SCNN(Pan et al., 2018) extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition.
+SCNN(Pan et al., 2018) modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer.
 RESA(Zheng et al., 2021) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information.
 Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
-HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
+HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
 2.2 Curve-based Lane Detection
 Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations.
-PolyLaneNet(Tabelini et al., 2021b) introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane.
-LSTR(Liu et al., 2021b) can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information.
+PolyLaneNet(Tabelini et al., 2021b) proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane.
+Similarly, LSTR(Liu et al., 2021b) utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions.
 DBNet(Dai et al., 2024) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation.
@@
 However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
+Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(GoogleMaps, Accessed: 2024)
+Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(GoogleMaps, Accessed: 2024)
 2.4 Row-wise-based Lane Detection
 Row-wise methods predict lane positions for each image row.
-E2E-LMD(Yoo et al., 2020) first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions.
+E2E-LMD(Yoo et al., 2020) introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly.
 UFLD(Qin et al., 2020) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes.
@@
 However, there is currently no publicly available foggy weather lane detection dataset.
-To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane.
+To fill this gap, we created the FoggyLane dataset for lane detection in foggy weather.
 In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017) datasets to test the generalizability of our methods.
-Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(GoogleMaps, Accessed: 2024)
-Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(GoogleMaps, Accessed: 2024)
+Figure 4: The different road types of FoggyLane images examples.
+Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
+Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.5k 1640 x 590
 3.1 FoggyLane
@@
 The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad.
-We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
-Figure 4: The different road types of FoggyLane images examples.
-Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
-Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
-Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
+The dataset was partitioned into training, validation and testing sets, with stratification to maintain a consistent distribution of scene types across both subsets.
+Com and paring FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
 While ApolloScape (Huang et al., 2018) and BDD100k (Yu et al., 2020) include multiple traffic scenarios, they lack instance-level annotations.
@@
 Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced.
-To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple (Tusimple, 2017) and CULane (Pan et al., 2018) formats for broader accessibility.
+To fill this gap, FoggyLane offers a thorough solution for lane detection in foggy conditions with instance-level annotations.
 Our dataset has two key advantages:
@@
 These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
+Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
+Figure 6: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
 3.2 FoggyCULane and FoggyTusimple
@@
 Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017).
-The CULane dataset (Pan et al., 2018) is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing.
+The CULane dataset (Pan et al., 2018) serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing.
 In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
 Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
-Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
 The method for generating foggy images is inspired by the model (Nie et al., 2022), and the complete generation process is shown in Fig.
@@
 The mathematical modeling equation defined by Koschmieder (Harald, 1924) is represented as:
-$I(x)=J(x)t(x)+A(1-t(x))$ (1)
+I(x)=J(x)t(x)+A(1t(x))I(x)=J(x)t(x)+A(1-t(x)) (1)
 where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image.
@@
 As this distance increases, the transmittance decreases, resulting in thicker fog.
-Figure 6: The FoggyCULane dataset examples.
 To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications.
@@
 Then they suppress information irrelevant to depth and use structural information for depth prediction.
-The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information.
-The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image.
-After passing through the DSA Module and STE Module, as shown in Fig.
-5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map.
-The Depth Prediction (DP) is responsible for depth estimation.
+The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details.
+The Depth-specific Attention (DSA) generates an attention map from the original image.
+Once processed through the DSA and STE Modules, as depicted in Fig.
+5, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map.
+The Depth Prediction (DP) module is then responsible for estimating the depth.
 Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
@@
 As illustrated in the Fig.
-8, the network comprises three main components: a backbone network, a neck network, and a detection head.
+8, the network comprises four main components.
 To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network.
@@
 4.2 Global Feature Fusion Module
-In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision.
+In recent years, Transformer architectures have made considerable strides in the field of natural language processing, leading researchers to explore their applications in computer vision.
 With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information.
-As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025b; Chen et al., 2023; Zhang et al., 2024b; Yang et al., 2023b; Zhang et al., 2025a).
+As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025; Chen et al., 2023).
 However, applying Transformers to the visual domain still faces some challenges.
@@
 Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
+Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
+[…]
+  • 1.
+    Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
+* indicates the FPS after acceleration with TensorRT.
 4.5 Loss functions
@@
 The Focal Loss between the predicted heatmap and the ground truth heatmap is defined as:
-$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$ (5)
+$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$
 4.5.2 Row-based segmentation prediction
@@
 FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach.
-It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing.
+It maintains the same number of images as CULane.
 The resolution of all images is also 1640x590 pixels.
-FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method.
+FoggyTusimple is synthesized from Tusimple through the foggy image modeling method.
 At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases.
 All visual data maintains a consistent resolution of 1280×720 pixels.
-Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
-  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN(Pan et al., 2018) AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.6 147.6 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.6 70.0 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.0 171.6 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.0 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(4.63$\uparrow$) 77.32(5.17$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.32$\uparrow$) 20.3 38.2 / 192.9*
-  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
-* indicates the FPS after acceleration with TensorRT.
 Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set.
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
+[…]
+  • 1.
+    Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow and Night conditions, as well as in the overall (Total) performance.
 5.2 Evaluation Metric
-Lane detection is commonly treated as a semantic segmentation task, evaluated using standard metrics such as True Positive (TP), False Positive (FP), False Negative (FN), and True Negative (TN).
-However, due to its unique structure, lane detection also adopts specialized metrics.
-Following SCNN’s(Pan et al., 2018) protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas.
-A prediction is considered correct if the Intersection over Union (IoU) with the ground truth exceeds 0.5.
-Based on this, we compute Precision, Recall, and F1-score, with the following formulas:
+Lane detection is typically approached as a semantic segmentation problem, assessed using common metrics.
+However, due to its distinct characteristics, lane detection also requires specialized evaluation metrics.
+Adhering to SCNN’s (Pan et al., 2018) framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions.
+Subsequently, we compute the Intersection over Union (IoU), and if the IoU value exceeds a predefined threshold, typically set at 0.5, the prediction is considered correct.
+From this, we calculate Precision, Recall, and F1-score using the following formulas:
 $F1=\frac{2\times Precision\times Recall}{Precision+Recall}$ (11)
@@
 $Recall=\frac{TP}{TP+FN}$ (13)
-For the FoggyTusimple dataset, we adopt the evaluation metrics of the Tusimple dataset.
-The official Tusimple dataset provides three standard evaluation metrics: false positive rate (FPR), false negative rate (FNR), and accuracy.
+Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+For the FoggyTusimple dataset, we use the same evaluation metrics as the Tusimple dataset.
+The official Tusimple dataset includes three standard metrics: false positive rate (FPR), false negative rate (FNR), and accuracy.
 The formula for calculating accuracy is as follows:
@@
 where $C_{\textit{clip}}$ represents the number of correctly predicted lane points in the image, while $S_{\textit{clip}}$ denotes the total number of ground truth lane points.
-For the pixels in the image, a predicted lane point is considered correct if it falls within 20 pixels to the left or right of the ground truth.
-Furthermore, if more than 85% of the predicted lane points on a single lane line correspond to the ground truth lane points, the predicted lane is regarded as a true positive.
+Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
+[…]
 5.3 Implementation Details
@@
 Various data augmentation techniques, including rotation, flipping, brightness adjustment, and random scaling, were applied.
-The batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay.
 For full network parameters and training details, please refer to our open-source code.
@@
 For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
-Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
-  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN(Pan et al., 2018) AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.07$\uparrow$) 78.63(0.87$\uparrow$) 73.89 82.74(1.85$\uparrow$) 53.75(0.50$\uparrow$) 89.82(0.37$\uparrow$) 74.09 1146 74.97(0.76$\uparrow$) 79.85(0.72$\uparrow$)
-  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow and Night conditions, as well as in the overall (Total) performance.
-Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
 5.4 Quantitative Results
@@
 We compared our method with other advanced lane detection methods on the FoggyCULane dataset.
-As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow and Curve.
+As shown in the Table 3, our method consistently achieved the best performance across various scenarios.
 Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
@@
 This further demonstrates the effectiveness of our method in foggy conditions.
-Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
-    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN(Pan et al., 2018) AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.89$\downarrow$) 3.83
 Figure 13: Several metrics of our model’s performance during the training process on the FoggyLane and FoggyTusimple datasets, with results on FoggyLane on the left and FoggyTusimple on the right.
@@
 92.06 11.97
-Ours 92.34$\uparrow$ 12.01
+[…]
 5.6 Model deployment
@@
 78.52 11.97
-Ours 78.66$\uparrow$ 12.01
+[…]
 Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin.
@@
 Furthermore, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, further validating its practical applicability in real-world scenarios.
-Declaration of competing interest
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
 Acknowledgement
@@
 2023B1212060029).We would like to release our source code and dataset.
-Author biography
-{biography}
-[zrh.jpg]Ronghui Zhang received a B.Sc.
-(Eng.) from the Department of Automation Science and Electrical Engineering, Hebei University, Baoding, China, in 2003, an M.S.
-degree in Vehicle Application Engineering from Jilin University, Changchun, China, in 2006, and a Ph.D.
-(Eng.) in Mechanical & Electrical Engineering from Changchun Institute of Optics, Fine Mechanics and Physics, the Chinese Academy of Sciences, Changchun, China, in 2009.
-After finishing his post-doctoral research work at INRIA, Paris, France, in February 2011, he is currently an Associate Professor with the Research Center of Intelligent Transportation Systems, School of intelligent systems engineering, Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China.
-His current research interests include computer vision, intelligent control and ITS.
-{biography}
-[mayh.jpg]Yuhang Ma received his B.S.
-degree in Automation at Huazhong University of Science and Technology in 2020.
-He is currently pursuing the master’s degree in Electronic and Information Engineering at Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China.
-His current research interests include autonomous driving and computer vision.
-{biography}
-[ltf.jpg]Tengfei Li received his B.S.
-degree in Robot Engineering from Harbin Institute of Technology at Weihai in 2024.
-He is currently pursuing the M.S.
-degree with the School of Intelligent Systems Engineering, Sun Yat-sen University.
-His research interests include computer vision and autonomous driving.
-{biography}[linziyu.pdf]Ziyu Lin is currently pursuing a B.Sc.
-degree in Traffic Engineering with the School of intelligent systems engineering, Sun Yat-sen University, Guangzhou, China.
-Her research interests include computer vision, deep learning, and autonomous driving technology, with a particular focus on exploring how these technologies can enhance the efficiency and safety of autonomous driving.
-{biography}
-[wyy.jpg]Yueying Wu received her B.S.
-degree in Internet of Things Engineering from Central South University, Changsha, China, in 2020, and an M.S.
-degree in Transportation Engineering from Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China, in 2023.
-Her current research interests include autonomous driving and vehicle-road collaboration, deep learning and computer vision.
-{biography}[cjz.png]Junzhou Chen received his Ph.D.
-in Computer Science and Engineering from the Chinese University of Hong Kong in 2008, following his M.Eng degree in Software Engineering and B.S.
-in Computer Science and Applications from Sichuan University in 2005 and 2002, respectively.
-Between March 2009 and February 2019, he served as a Lecturer and later as an Associate Professor at the School of Information Science and Technology at Southwest Jiaotong University.
-He is currently an associate professor at the School of Intelligent Systems Engineering at Sun Yat-sen University.
-His research interests include computer vision, machine learning, intelligent transportation systems, mobile computing and medical image processing.
-{biography}
-[…]

### v10

@@
 This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments.
-To address this, we introduce the FoggyLane dataset, captured in real-world foggy scenarios, and synthesize two additional datasets, FoggyCULane and FoggyTusimple, from existing popular lane detection datasets.
+To address this, we introduce the FoggyLane dataset, a diverse and comprehensive benchmark captured in real-world foggy scenarios across multiple road types and conditions, and synthesize two more datasets, FoggyCULane and FoggyTusimple, derived from widely used lane detection datasets.
 Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions.
@@
 ADAS, Intelligent vehicle, Foggy lane detection dataset, Foggy lane detection, Global attention, Edge enhancement
-journal: Expert Systems With Applications
+journal: Neurocomputing
 […]
@@
 [5]organization=State Key Laboratory of Automotive Chassis Integration and Bionics, Jilin University, city=Changchun, postcode=130025, country=China
+{highlights}
+A new real-world dataset and two synthetic datasets for foggy lane detection are proposed.
+A novel Fog-Enhanced Network is proposed for lane detection in fog.
+Three core modules are designed to enhance feature extraction in fog.
+The method achieves SOTA performance with real-time speed on embedded devices.
 1 Introduction
-Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents.
-As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Chen et al., 2026).
+Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents[1].
+As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving[2].
 Under good weather and clear lane markings, current lane detection algorithms perform well.
-However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging(Sang and Norris, 2025).
+However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging.
 In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences.
@@
 This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
-Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels, 2025)
-Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2).
-In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments.
-Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions.
-For example, Zhang and Zhu (2019) proposed an autonomous path tracking control method for intelligent electric vehicles based on lane detection and an optimal preview strategy, demonstrating that accurate lane recognition and trajectory tracking are fundamental to ensuring driving stability and safety at lower automation levels (Zhang and Zhu, 2019).
-Similarly, recent research highlights that coordinated lane-changing strategies in mixed traffic, such as the decentralized multi-vehicle decision-making framework developed by Alshakhs et al., are essential for improving both traffic efficiency and safety in environments where human-driven vehicles and connected automated vehicles coexist (Alshakhs et al., 2025).
-Therefore, this research not only focuses on end-to-end autonomous driving technologies but also aims to improve lane detection performance in complex environments, providing stronger safety support and driving assistance for L0, L1, and L2 vehicles in adverse conditions.
-Earlier lane detection algorithms mainly relied on traditional computer vision methods.
-Son et al.(Son et al., 2015) first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering.
-Hou et al.(Hou et al., 2016) used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines.
-While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
-With deep learning’s success in many traffic tasks(Zhang et al., 2024b; Shunmuga Perumal et al., 2023; Preeti and Rana, 2024; Awan et al., 2025), it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization.
-Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018).
-However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation.
-Qin et al.(Qin et al., 2020) addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
-In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection(Redmon, 2016; Farhadi and Redmon, 2018).
-Due to the elongated shape of lane lines, traditional bounding boxes are unsuitable.
-In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position(Chen et al., 2019; Zheng et al., 2022; Ran et al., 2023; Wang et al., 2025).
-These innovations improve accuracy and efficiency in complex or occluded scenarios.
-Research on downstream tasks in adverse weather conditions, such as fog, is limited compared to other detection and segmentation tasks.
-Traditional solutions often use image enhancement techniques, like dehazing, to improve visual clarity, but this doesn’t always lead to better detection results.
-Some methods combine image enhancement with detection to reduce weather interference(Huang et al., 2020), while others simulate foggy images to augment datasets.
-For example, Nie et al.(Nie et al., 2022) generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
-Domain adaptation techniques also address the style gap between foggy and clear-weather images(Chen et al., 2018; Zhang et al., 2021).
-Hnewa et al.(Hnewa and Radha, 2021) proposed a multi-scale domain-adaptive YOLO model, while Li et al.(Li et al., 2023) introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
-Traditional dehazing methods suffer from artifacts, whereas deep learning methods using CNNs capture more complex features but require high computational resources and diverse training datasets.
-However, obtaining real paired images under foggy conditions is challenging, and synthetic images often lead to reduced performance in real-world scenarios.
+Figure 1: Lane detection in foggy scenarios for ADAS.[3]
+Earlier lane detection algorithms [4, 5] primarily relied on traditional computer vision methods, such as feature-based approaches that detect vanishing points, use color features, and apply edge detection or line detection techniques like Hough transforms.
+These methods are efficient for real-time applications but require manual parameter tuning and often struggle in challenging scenarios like worn or occluded lane markings.
+With the rise of deep learning, many modern lane detection methods [6, 7, 8, 9, 10, 11] leverage CNNs, which automatically extract features from raw data and generalize better in complex road environments.
+While deep learning provides higher robustness, many models still rely on pixel-by-pixel classification, which is computationally expensive and unnecessary for lane detection.
+Moreover, lane detection doesn’t always require dense pixel segmentation, which can be inefficient.
+Additionally, some approaches have incorporated anchor-based methods from object detection to improve efficiency and accuracy, particularly for complex or occluded lanes.
+These methods preset anchors at different starting points and slopes to optimize lane shape and position prediction [12, 13, 14, 15, 16].
+However, these techniques still face challenges in handling foggy conditions, where image degradation severely impacts lane visibility.
+Traditional image enhancement methods, like dehazing, are commonly used to improve clarity but do not always result in better detection performance [17].
+Some approaches generate synthetic foggy images for data augmentation, but these do not always capture real-world complexities [18].
+Moreover, domain adaptation techniques [19, 20, 21, 22, 23] have been employed to bridge the gap between foggy and clear-weather images.
+These methods often rely on labeled clear-weather datasets or use transfer learning, but they do not fully address the unique challenges posed by foggy conditions.
 Lane detection enables key functionalities such as lane departure warnings and autonomous navigation in ADAS.
@@
 This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets.
-Additionally, we create foggy versions of two commonly used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), resulting in the FoggyCULane and FoggyTusimple datasets.
+Additionally, we create foggy versions of two commonly used lane detection datasets, CULane [8] and Tusimple [24], resulting in the FoggyCULane and FoggyTusimple datasets.
 These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@
   • 3)
-    To tackle the loss of edge information in foggy images, we incorporate a Low-level Edge Enhancement Module (LEEM) that enhances the model’s attention to edge features, improving the accuracy and stability of lane detection.
+    To tackle the loss of edge information in foggy images, we incorporate a Low-level Edge Enhancement Module (LEEM), which improves the model’s sensitivity to fine edge details, ensuring better lane delineation even when visibility is severely impaired by fog.
   • 4)
@@
 Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem.
-LaneNet(Neven et al., 2018) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines.
-SCNN(Pan et al., 2018) modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer.
-RESA(Zheng et al., 2021) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information.
+LaneNet[6] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines.
+SCNN[8] modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer.
+RESA[25] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information.
 Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
-HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
+HW-Transformer[26] presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
 2.2 Curve-based Lane Detection
 Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations.
-PolyLaneNet(Tabelini et al., 2021b) proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane.
-Similarly, LSTR(Liu et al., 2021b) utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions.
-DBNet(Dai et al., 2024) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation.
+PolyLaneNet[27] proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane.
+Similarly, LSTR[28] utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions.
+DBNet[29] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation.
 However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
@@
 Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings.
-Line-CNN(Li et al., 2019) introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors.
-LaneATT(Tabelini et al., 2021a) effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module.
-CLRNet(Zheng et al., 2022) employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features.
+Line-CNN[30] introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors.
+LaneATT[31] effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module.
+CLRNet[16] employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features.
 This hierarchical fusion of semantic information significantly improves detection precision.
-FLAMNet(Ran et al., 2023) employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms.
+FLAMNet[32] employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms.
 This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures.
 However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
-Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(GoogleMaps, Accessed: 2024)
-Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(GoogleMaps, Accessed: 2024)
+Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.[33]
+Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[33]
 2.4 Row-wise-based Lane Detection
 Row-wise methods predict lane positions for each image row.
-E2E-LMD(Yoo et al., 2020) introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly.
-UFLD(Qin et al., 2020) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes.
-CondLaneNet(Liu et al., 2021a) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts.
+E2E-LMD[34] introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly.
+UFLD[35] utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes.
+CondLaneNet[36] utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts.
 The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
@@
 To fill this gap, we created the FoggyLane dataset for lane detection in foggy weather.
-In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017) datasets to test the generalizability of our methods.
+In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [8] and Tusimple [24] datasets to test the generalizability of our methods.
 Figure 4: The different road types of FoggyLane images examples.
 Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
-Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.5k 1640 x 590
+Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [37] 1.2k None 640 x 480 Tusimple [24] 6.4k None 1280 x 720 LLAMAS [38] 100k None 1276 x 717 Appolloscape [39] 110k None 3384 x 2710 BDD100k [40] 100k None 1280 x 720 CULane [8] 133k None 1640 x 590 CurveLanes [41] 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.423k 1640 x 590
 3.1 FoggyLane
 Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents.
-Current benchmark datasets, including CULane(Pan et al., 2018), Tusimple(Tusimple, 2017), and LLAMAS(Behrendt and Soussan, 2019), lack foggy weather data.
+Current benchmark datasets, including CULane[8], Tusimple[24], and LLAMAS[38], lack foggy weather data.
 To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions.
@@
 The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
-Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube, Accessed: 2024), resulting in an additional 337 images.
+Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos[42], resulting in an additional 337 images.
 The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad.
 The dataset was partitioned into training, validation and testing sets, with stratification to maintain a consistent distribution of scene types across both subsets.
-Com and paring FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
-While ApolloScape (Huang et al., 2018) and BDD100k (Yu et al., 2020) include multiple traffic scenarios, they lack instance-level annotations.
-Similarly, although CULane (Pan et al., 2018) and CurveLanes (Xu et al., 2020) cover a wide range of traffic scenes, they do not include foggy conditions.
+Comparing FoggyLane with existing datasets, such as Caltech-Lanes[37] and Tusimple[24], which focus on single traffic scenarios, and LLAMAS[38], which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
+While ApolloScape [39] and BDD100k [40] include multiple traffic scenarios, they lack instance-level annotations.
+Similarly, although CULane [8] and CurveLanes [41] cover a wide range of traffic scenes, they do not include foggy conditions.
 Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced.
@@
 Our dataset has two key advantages:
-  1. 1.
+  • 1)
     Pioneering foggy lane detection dataset: FoggyLane is a dataset specifically designed to address foggy weather scenarios.
@@
 This dataset is crucial for testing the robustness of ADAS under adverse conditions.
-  2. 2.
-    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.(Zhang et al., 2024a).
+  • 2)
+    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[43].
 These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
@@
 Currently, numerous datasets are available for lane detection.
-Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017).
-The CULane dataset (Pan et al., 2018) serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing.
-In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
+Among them, we select two widely adopted and representative benchmarks: CULane[8] and TuSimple[24].
+The CULane dataset [8] serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing.
+In contrast, the TuSimple dataset [24] primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
 Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
-The method for generating foggy images is inspired by the model (Nie et al., 2022), and the complete generation process is shown in Fig.
+The method for generating foggy images is inspired by the model [18], and the complete generation process is shown in Fig.
 5.
@@
 It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one.
-The mathematical modeling equation defined by Koschmieder (Harald, 1924) is represented as:
+The mathematical modeling equation defined by Koschmieder [44] is represented as:
 I(x)=J(x)t(x)+A(1t(x))I(x)=J(x)t(x)+A(1-t(x)) (1)
@@
 As this distance increases, the transmittance decreases, resulting in thicker fog.
-To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications.
+To calculate the corresponding $d(x)$, we use the S2R-DepthNet[45] that has good generalization performance across various applications.
 The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image.
@@
 According to the transmittance formula Eq.
-(2), we use the dark channel prior(He et al., 2011) to calculate the average atmospheric light value $A$.
+(2), we use the dark channel prior[46] to calculate the average atmospheric light value $A$.
 In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel.
@@
 4.1 Overall structure
-To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods.
-As illustrated in the Fig.
-8, the network comprises four main components.
-To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network.
-In the neck network, we utilize HRNetV2(Wang et al., 2021) to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
+To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network.
+The network works through several collaborative modules to tackle the challenges posed by foggy images, with the overall architecture shown in Fig.
+8.
+The input image first passes through the backbone network, where feature extraction occurs, followed by the Global Feature Fusion Module (GFFM).
+This module is designed to mitigate the loss of global information in foggy images, allowing the network to retain essential global context.
+The output is a feature map enriched with both local and global information, crucial for accurate lane detection.
+The multi-level feature map from the backbone is then passed into the Neck section of the network, which employs HRNetV2[47] for multi-scale feature fusion.
+HRNetV2 excels in integrating high-resolution and low-resolution features from different stages of the backbone.
+This process helps the network capture spatial information at multiple scales, producing a rich, multi-scale feature representation that forms the basis for lane instance detection.
+Next, the Instance Detection Head processes the multi-scale feature map from the Neck network, predicting dynamic convolution kernel parameters for each lane instance.
+These dynamic kernels are applied to the lane mask segmentation head to isolate instance-specific features.
+The output includes a parameter map and a heat map, crucial for identifying lane instances.
+This approach eliminates the need for complex post-processing, significantly enhancing detection speed and accuracy.
+The output parameter map is then passed to the Kernel Feature Fusion Module (KFFM), which learns the relationships between different lane instances.
+The KFFM fuses these instance-specific features, improving lane localization accuracy, especially for complex or overlapping lane lines.
+This module ensures that the network can better identify lane instances and refine the localization of each lane.
+Finally, to compensate for the loss of edge information in foggy images, the network incorporates a Low-Level Edge Enhancement Module (LEEM).
+This module enhances edge features using convolution operations, recovering fine details at lane boundaries.
+The enhanced edge features improve the network’s ability to detect lane boundaries with high precision, even under challenging foggy conditions.
 Figure 8: Overview of the proposed network.
-For the detection head, we employ the concept of dynamic convolution, incorporating both lane instance detection and lane mask segmentation.
-During instance detection, the network identifies lane instances within the image and predicts a set of dynamic convolution kernel parameters for each lane.
-These parameters are subsequently applied to the lane mask segmentation head to filter the instance-specific segmentation map for each lane.
-This approach eliminates the need for complex post-processing steps, enhancing detection speed.
-Moreover, to fully exploit the structural relationships between lane lines, we designed a kernel feature fusion module, which allows the network to automatically learn the correlations between predicted lane instances.
-This contributes to more accurate lane identification and localization.
-Additionally, to compensate for the loss of edge information in foggy images, we incorporated a low-level edge enhancement module, further improving the network’s ability to detect lane boundaries with precision.
 4.2 Global Feature Fusion Module
-In recent years, Transformer architectures have made considerable strides in the field of natural language processing, leading researchers to explore their applications in computer vision.
-With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information.
-As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025; Chen et al., 2023).
-However, applying Transformers to the visual domain still faces some challenges.
-Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities.
-Additionally, calculating global attention at high resolutions can be computationally intensive.
+As shown in Fig.
+9, to enhance global feature integration in the Swin Transformer backbone[48], we replace the final window-based attention mechanism with a standard multi-head attention mechanism (MHSA) from Vision Transformer (ViT)[49].
+While ViT excels at capturing global dependencies, it suffers from high computational cost, especially at high resolutions, making it unsuitable for real-time tasks such as lane detection.
+The traditional ViT computes global self-attention with a complexity of $O(N^{2})$, where $N$ is the number of input tokens.
+This high computational cost becomes prohibitive at high resolution, slowing down inference and limiting real-time processing.
+In our approach, we address this issue by introducing an efficient solution: global attention is applied only after reducing the image resolution.
+This ensures that the computational cost of global attention is manageable, allowing us to retain the global attention mechanism of ViT while maintaining fast inference speed, which is crucial for real-time lane detection in foggy conditions.
+In our design, this MHSA layer replaces the original window-based attention mechanism used in the Swin Transformer.
+By applying global attention after reducing the image resolution, we strike a balance between maintaining global context awareness and achieving high computational efficiency, which is crucial for real-time lane detection tasks.
 Figure 9: Backbone Network with Global Feature Fusion.
-To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers(Liu et al., 2021c, 2022; Dong et al., 2022).
-Among them, the Swin Transformer(Liu et al., 2021c) combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework.
-As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network.
-However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding.
-Therefore, as illustrated in the Fig.
-9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration.
-Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs.
-At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
 4.3 Kernel Feature Fusion Module
-Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution.
-CondInst(Tian et al., 2023) is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency.
-Similarly, CondLane(Liu et al., 2021a) applies dynamic convolution for lane instance segmentation, achieving strong performance.
-Building on this, our lane detection head also uses dynamic convolution, consisting of lane instance detection and mask segmentation.
-Unlike traditional methods, we generate dynamic convolution kernels using the lane’s starting position features rather than center features, which are more suitable for the elongated nature of lanes.
-We design an instance detection head with two branches: one predicts lane starting points and generates a heatmap, and the other generates dynamic convolution kernels for lane segmentation.
-However, existing methods don’t fully utilize the structural relationships between lane instances.
-To address this, we introduce a Kernel Feature Fusion Module (KFFM) that enables the network to learn and predict relationships between lane instances, enhancing detection performance.
-KFFM is shown in the bottom right of Fig.
-8.
+Existing lane instance segmentation methods[50, 36], typically use dynamic convolution to generate kernel parameters for each instance, which improves the network’s flexibility and adaptability.
+However, these methods do not fully leverage the structural relationships between lane instances, leading to limited segmentation and detection accuracy.
+Traditional dynamic convolution methods only consider individual lane instances and overlook the potential relationships between adjacent instances, which reduces their performance when handling complex lane scenes.
+To address this limitation, we propose the Kernel Feature Fusion Module (KFFM), as shown in the bottom right of Fig.
+8, which optimizes kernel generation by integrating information from adjacent lane instances.
+Specifically, KFFM introduces structured feature fusion into dynamic convolution, allowing each lane instance’s kernel to depend not only on the features of the current lane but also on the information from nearby lane instances.
+This approach enables the network to adaptively learn the relationships between lane instances, improving both detection accuracy and segmentation performance.
+Let the input feature map from the lane instance detection branch be denoted as $\mathit{K}=\left[\mathit{kernel}_{0},\mathit{kernel}_{1},\dots,\mathit{kernel}_{N}\right]$, where each $\mathit{kernel}_{i}\in\mathbb{R}^{k\times k\times C}$, and define the weight matrix $\mathit{W}\in\mathbb{R}^{(N+1)\times(N+1)}$ to capture the contributions of each kernel to the final fused kernel.
+The final fused kernel for lane $i$ is computed as the weighted sum of the kernels:
+$\hat{\mathit{kernel}}_{i}=\sum_{j=0}^{N}W_{i,j}\cdot\mathit{kernel}_{j}$ (4)
+where $W_{i,j}$ is the weight for the contribution of $\mathit{kernel}_{j}$ to the fused kernel $\hat{\mathit{kernel}}_{i}$; The weight matrix $\mathit{W}$ is learned during training.
+The final fused kernel is obtained through matrix multiplication as follows:
+$\hat{\mathit{K}}_{i}=\mathit{W}_{i}\cdot\mathit{K}$ (5)
+where $\mathit{W}_{i}$ corresponds to the row of the weight matrix $\mathit{W}$ associated with kernel $i$.
+The resulting fused kernel $\hat{\mathit{kernel}}_{i}$ is then used in the dynamic convolution operation for lane segmentation, leading to more accurate lane detection and segmentation.
 4.4 Low-level Edge Enhanced Module
-Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast.
-This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately.
-Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking.
-Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig.
+Images captured in foggy conditions differ from those captured in good lighting primarily due to blurriness and low contrast, which makes the edges of objects and scene details indistinct and difficult to detect.
+This degradation of edge information significantly impacts the algorithm’s ability to accurately understand and recognize the scene.
+Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial relationships about objects, which is critical for tasks such as object detection, segmentation, and tracking.
+We used the Canny edge detection algorithm with identical threshold settings to process both foggy and normal images, as shown in Fig.
 10.
-It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible.
-Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
+The results show that fog significantly degrades edge information, particularly lane line edges, which become nearly indiscernible.
+Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—accurate edge detection is essential for reliable lane detection.
+To address the challenges posed by foggy conditions, we propose the integration of global feature fusion within both the backbone network and detection head to enhance high-dimensional global features.
+Additionally, we incorporate multi-scale fusion to aggregate detailed features from low-resolution feature maps.
+However, this approach alone does not fully compensate for the loss of edge details in foggy images.
+To resolve this issue, we introduce the Low-level Edge Enhanced Module (LEEM) into the network.
+LEEM focuses on edge information, effectively recovering and refining the edges that are crucial for lane detection, thus improving both the accuracy and stability of lane detection in foggy environments.
 Figure 10: Comparison of edge information between foggy and normal scenarios.
-To address the challenges of lane detection in foggy conditions, we apply global feature fusion within both the backbone network and detection head to enhance high-dimensional global features.
-Additionally, we use multi-scale fusion to aggregate detail features from low-resolution feature maps.
-However, this approach alone does not fully compensate for the loss of edge detail inherent in foggy images.
-Therefore, we introduce a Low-level Edge Enhanced Module (LEEM) into the network to boost the network’s focus on edge information, thereby improving both the accuracy and stability of lane detection.
-Specifically, we add an auxiliary path in the network to learn the edge features of the input image.
-First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module.
-To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network.
-This branch sequentially applies a 3 × 3 and a 1 × 1 convolution.
-We then compute the cross-entropy loss between the branch output and the edge labels.
-The detailed structure is illustrated in the bottom-left of the Fig.
-8.
-Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
+LEEM works by adding an auxiliary path to the network that specifically learns edge features of the input image.
+The process begins with the application of the Canny operator for edge detection, followed by the reconstruction of missing lane edges using lane line labels.
+This ground truth for the edge enhancement module is generated by combining detected edges with labeled edges.
+This approach helps the network better identify lane boundaries, which are often difficult to distinguish in foggy conditions.
+In the second stage of the backbone network, we introduce a branch that processes the extracted edge features.
+This branch sequentially applies a $3\times 3$ convolution followed by a $1\times 1$ convolution to enhance edge-related features.
+The mathematical representation of this process is as follows:
+$\hat{\mathit{E}}_{\mathit{predict}}=\mathit{Conv}_{2}\left(\mathit{ReLU}\left(\mathit{BN}(\mathit{Conv}_{1}(\mathit{X})))\right)\right)$ (6)
+where $\mathit{Conv}_{1}$ is a $3\times 3$ convolution designed to capture spatial edge features, $\mathit{BN}$ is the batch normalization layer to stabilize the learning process by normalizing the output, $\mathit{ReLU}$ is the activation function that provides non-linearity, $\mathit{Conv}_{2}$ is a $1\times 1$ convolution used to reduce the dimensionality and refine edge features.
+During training, the edge information extracted via the Canny operator from both the input image and ground truth (GT) lane lines is first downsampled to match the resolution of the network’s predicted edge information $\hat{\mathit{E}}_{\mathit{predict}}$.
+Then, these two edge information maps are merged.
+The final combined edge information is represented as $\hat{\mathit{E}}_{\mathit{GT}}$, which can be computed as:
+$\hat{\mathit{E}}_{\mathit{GT}}=\mathit{Edge}_{I}+\mathit{Edge}_{\mathit{Lane}}$ (7)
+where $\mathit{Edge}_{I}$ is the edge information obtained from the input image $\mathit{I}$ using the Canny operator, $\mathit{Edge}_{\mathit{Lane}}$ is the edge information derived from the lane ground truth (GT) using the Canny operator.
+Next, we compute the cross-entropy loss between the predicted edge information $\hat{\mathit{E}}_{\mathit{predict}}$ and the final combined edge information $\hat{\mathit{E}}_{\mathit{GT}}$.
+This loss function is formulated as:
+$L_{\mathit{edge}}=-\sum_{i}\left(e_{\mathit{gt}}^{i}\log(p_{i})+(1-e_{\mathit{gt}}^{i})\log(1-p_{i})\right)$ (8)
+where $e_{\mathit{gt}}^{i}$ represents the ground truth label for pixel $i$ (1 if it’s part of the edge, 0 otherwise), $p_{i}$ is the predicted probability that pixel $i$ is part of the edge.
+The LEEM module allows the network to effectively enhance critical edge details in lane detection, significantly improving both the accuracy and stability of lane detection in foggy conditions.
+This module not only recovers lost edge information in images but also integrates information from various sources, using the cross-entropy loss function to optimize the network.
+As a result, LEEM plays a key role in enhancing the robustness of lane detection systems in challenging environments.
 Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
@@
 * indicates the FPS after acceleration with TensorRT.
+Figure 11: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyLane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
+[…]
+  • 1.
+    Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow and Night conditions, as well as in the overall (Total) performance.
 4.5 Loss functions
@@
 Accordingly, the cross-entropy loss function is employed for optimization.
-Let $e_{gt}^{i}$ denote the label of pixel $i$, and $p_{i}$ represent the predicted probability of the pixel being positive.
-The loss function is defined as:
+As discussed in Section 4.4, the loss function for edge prediction is defined as:
 $L_{\text{edge}}=\sum_{i}\left(-e_{gt}^{i}\log(p_{i})-(1-e_{gt}^{i})\log(1-p_{i})\right)$ (14)
+where $e_{gt}^{i}$ denotes the label of pixel $i$, and $p_{i}$ represents the predicted probability of the pixel being part of the edge.
 4.5.6 Total loss
@@
 All visual data maintains a consistent resolution of 1280×720 pixels.
-Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
-[…]
-  • 1.
-    Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow and Night conditions, as well as in the overall (Total) performance.
 5.2 Evaluation Metric
@@
 However, due to its distinct characteristics, lane detection also requires specialized evaluation metrics.
-Adhering to SCNN’s (Pan et al., 2018) framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions.
+Adhering to SCNN’s [8] framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions.
 Subsequently, we compute the Intersection over Union (IoU), and if the IoU value exceeds a predefined threshold, typically set at 0.5, the prediction is considered correct.
@@
 $Recall=\frac{TP}{TP+FN}$ (18)
-Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set.
+Figure 12: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyCULane testing set.
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@
 For full network parameters and training details, please refer to our open-source code.
-To benchmark against existing methods(Pan et al., 2018; Zheng et al., 2021; Tabelini et al., 2021a; Zheng et al., 2022; Ran et al., 2023; Liu et al., 2021b; Dai et al., 2024; Qin et al., 2020; Liu et al., 2021a; Wang et al., 2025), we followed the training settings provided in the respective papers.
+To benchmark against existing methods[8, 35, 36, 25, 31, 28, 16, 32, 29, 51], we followed the training settings provided in the respective papers.
 For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
@@
 2022B1515120072), the Guangzhou Science and Technology Plan Project (No.
-2024B01W0079), the Nansha Key RD Program (No.
-2022ZD014), the Science and Technology Planning Project of Guangdong Province (No.
+2024B01W0079), the Science and Technology Planning Project of Guangdong Province (No.
 2023B1212060029).We would like to release our source code and dataset.

### v11

@@
-A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions
+
 Abstract
-Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS).
-Existing lane detection algorithms generally perform well under favorable weather conditions.
-However, their performance degrades significantly in adverse conditions, such as fog, which increases the risk of traffic accidents.
-This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments.
-To address this, we introduce the FoggyLane dataset, a diverse and comprehensive benchmark captured in real-world foggy scenarios across multiple road types and conditions, and synthesize two more datasets, FoggyCULane and FoggyTusimple, derived from widely used lane detection datasets.
-Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions.
-Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, with F1-scores of 95.04 on FoggyLane, 79.85 on FoggyCULane, and 96.95 on FoggyTusimple.
-Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
+Lane departure accident prevention plays a critical role in enhancing road safety, and lane detection is a core technology to achieve this goal, especially under complex weather conditions.
+While existing lane detection algorithms perform well under favorable weather conditions, their effectiveness significantly degrades in foggy environments, which increases the risk of traffic accidents.
+In response to this challenge, we propose PDT-Net, a robust Prior-Guided Dynamic Feature Fusion Transformer framework designed for real-time lane detection in foggy conditions.
+This framework integrates three key modules: a Global Feature Fusion Module (GFFM) to capture the relationship between local and global features in foggy images, a Dynamic Feature Fusion Module (DFFM) to model the structural and positional relationships of lane instances, and a Prior-Guided Edge Enhancement Module (PEM) to recover lost edge details in foggy environments.
+Furthermore, we introduce the FoggyLane dataset, a real-world dataset that specifically targets lane detection in foggy conditions, along with two synthesized datasets, FoggyCULane and FoggyTusimple, to address the lack of fog-specific data for lane detection.
+Extensive experiments show that PDT-Net achieves state-of-the-art performance with F1-scores of 95.04% on FoggyLane, 79.85% on FoggyCULane, and 96.95% on FoggyTusimple.
+Moreover, with TensorRT acceleration, our method achieves a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capability and robustness in challenging foggy environments.
+By improving the precision of lane detection, our framework can contribute to active safety warning systems, helping to prevent accidents in foggy conditions.
 keywords:
-ADAS, Intelligent vehicle, Foggy lane detection dataset, Foggy lane detection, Global attention, Edge enhancement
-journal: Neurocomputing
 […]
-[1]organization=Guangdong Key Laboratory of Intelligent Transportation System, School of Intelligent Systems Engineering, Sun Yat-sen University, city=Guangzhou, postcode=510275, country=China
-[…]
-[2]organization=College of Automotive Studies, Tongji University, city=Shanghai, postcode=201804, country=China
-[…]
-[3]organization=College of Transportation Engineering, Tongji University, city=Shanghai, postcode=201804, country=China
-[…]
-[4]organization=Department of Civil and Environmental Engineering, University of Alberta, city=Edmonton, state=Alberta, country=Canada
-[…]
-[5]organization=State Key Laboratory of Automotive Chassis Integration and Bionics, Jilin University, city=Changchun, postcode=130025, country=China
-{highlights}
-A new real-world dataset and two synthetic datasets for foggy lane detection are proposed.
-A novel Fog-Enhanced Network is proposed for lane detection in fog.
-Three core modules are designed to enhance feature extraction in fog.
-The method achieves SOTA performance with real-time speed on embedded devices.
 1 Introduction
-Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents[1].
-As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving[2].
+Road traffic crashes continue to be a major global safety concern, with lane departure being a significant contributor.
+According to the National Highway Traffic Safety Administration (NHTSA), over 40% of fatal accidents in 2023 were attributed to lane departure(NHTSA2025), primarily caused by adverse weather conditions, degraded road markings, and driver errors.
+These factors expose critical gaps in vehicle perception, highlighting the need for more robust lane detection systems to address this issue.
+Computer vision techniques, particularly those focused on real-time risk assessment and accident prediction, play a pivotal role in enhancing active safety warning systems(AAP4).
+Lane detection, as a core perceptual capability, enables Advanced Driver Assistance Systems (ADAS) to trigger active safety warnings such as lane departure warning, which are essential for accident prevention.
 Under good weather and clear lane markings, current lane detection algorithms perform well.
@@
 As shown in Fig.
-1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module.
-Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud.
-This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
-Figure 1: Lane detection in foggy scenarios for ADAS.[3]
-Earlier lane detection algorithms [4, 5] primarily relied on traditional computer vision methods, such as feature-based approaches that detect vanishing points, use color features, and apply edge detection or line detection techniques like Hough transforms.
-These methods are efficient for real-time applications but require manual parameter tuning and often struggle in challenging scenarios like worn or occluded lane markings.
-With the rise of deep learning, many modern lane detection methods [6, 7, 8, 9, 10, 11] leverage CNNs, which automatically extract features from raw data and generalize better in complex road environments.
+1, The in-vehicle camera captures real-time road images, which are then preprocessed by edge devices and sent to the lane detection module.
+The detected lane information is used for active safety alerts.
+In the event of lane departure, the system proactively issues warnings(AAP7) through steering wheel vibration, voice notifications, and seat vibration, helping to prevent accidents(AAP8).
+Earlier lane detection algorithms(son2015real; hou2016efficient) primarily relied on traditional computer vision methods, such as feature-based approaches that detect vanishing points, use color features, and apply edge detection or line detection techniques like Hough transforms.
+These methods are efficient for real-time applications but require manual parameter tuning and often struggle in challenging scenarios like worn or occluded lane markings, which may lead to serious collision accidents.
+Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels) Inaccurate lane detection may lead to lane departure, increasing the risk of collision.
+Accurate lane detection provides essential technical support for the active safety functions of vehicles.
+Once the vehicle deviates from its intended lane, the ADAS can promptly issue an active safety warning to alert the driver, enabling timely corrective actions and ultimately preventing the occurrence of traffic accidents.
+With the rise of deep learning, many modern lane detection methods (neven2018towards; yu2018deep; pan2018spatial; abualsaud2021laneaf) leverage CNNs, which automatically extract features from raw data and generalize better in complex road environments.
 While deep learning provides higher robustness, many models still rely on pixel-by-pixel classification, which is computationally expensive and unnecessary for lane detection.
@@
 Additionally, some approaches have incorporated anchor-based methods from object detection to improve efficiency and accuracy, particularly for complex or occluded lanes.
-These methods preset anchors at different starting points and slopes to optimize lane shape and position prediction [12, 13, 14, 15, 16].
+These methods preset anchors at different starting points and slopes to optimize lane shape and position prediction (AAP1; redmon2016you; farhadi2018yolov3; chen2019pointlanenet; zheng2022clrnet; AAP5).
 However, these techniques still face challenges in handling foggy conditions, where image degradation severely impacts lane visibility.
-Traditional image enhancement methods, like dehazing, are commonly used to improve clarity but do not always result in better detection performance [17].
-Some approaches generate synthetic foggy images for data augmentation, but these do not always capture real-world complexities [18].
-Moreover, domain adaptation techniques [19, 20, 21, 22, 23] have been employed to bridge the gap between foggy and clear-weather images.
+Traditional image enhancement methods, like dehazing, are commonly used to improve clarity but do not always result in better detection performance (huang2020dsnet).
+Some approaches generate synthetic foggy images for data augmentation, but these do not always capture real-world complexities (nie2022sensor).
+Moreover, domain adaptation techniques (chen2018domain; zhang2021domain; hnewa2021multiscale; li2023domain) have been employed to bridge the gap between foggy and clear-weather images.
 These methods often rely on labeled clear-weather datasets or use transfer learning, but they do not fully address the unique challenges posed by foggy conditions.
-Lane detection enables key functionalities such as lane departure warnings and autonomous navigation in ADAS.
+Lane detection enables key functionalities such as lane departure warnings and autonomous navigation in ADAS, demonstrating strong risk perception capabilities(AAP3; AAP6).
 While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents.
@@
 This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets.
-Additionally, we create foggy versions of two commonly used lane detection datasets, CULane [8] and Tusimple [24], resulting in the FoggyCULane and FoggyTusimple datasets.
+Additionally, we create foggy versions of two commonly used lane detection datasets, CULane (pan2018spatial) and Tusimple (Tusimple), resulting in the FoggyCULane and FoggyTusimple datasets.
 These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@
   • 2)
-    To enhance the utilization of structural relationships between lane instances, we introduce a Kernel Feature Fusion Module (KFFM) that learns and predicts the correlations between lane instances, improving lane detection accuracy and robustness.
+    To enhance the utilization of structural relationships between lane instances, we introduce a Dynamic Feature Fusion Module (DFFM) that learns and predicts the correlations between lane instances, improving lane detection accuracy and robustness.
   • 3)
-    To tackle the loss of edge information in foggy images, we incorporate a Low-level Edge Enhancement Module (LEEM), which improves the model’s sensitivity to fine edge details, ensuring better lane delineation even when visibility is severely impaired by fog.
+    To tackle the loss of edge information in foggy images, we incorporate a Prior-Guided Edge Enhancement Module (PEM), which enhances the model’s sensitivity to fine edge details.
+This ensures better lane delineation, even when visibility is severely impaired by fog, and significantly improves the reliability of lane departure warnings.
   • 4)
@@
 Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem.
-LaneNet[6] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines.
-SCNN[8] modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer.
-RESA[25] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information.
+LaneNet(neven2018towards) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines.
+SCNN(pan2018spatial) modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer.
+RESA(zheng2021resa) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information.
 Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources.
-HW-Transformer[26] presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
+HW-Transformer(tits3) presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
 2.2 Curve-based Lane Detection
 Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations.
-PolyLaneNet[27] proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane.
-Similarly, LSTR[28] utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions.
-DBNet[29] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation.
+PolyLaneNet(tabelini2021polylanenet) proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane.
+Similarly, LSTR(liu2021end) utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions.
+DBNet(dai2024dbnet) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation.
 However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
@@
 Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings.
-Line-CNN[30] introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors.
-LaneATT[31] effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module.
-CLRNet[16] employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features.
+Line-CNN(li2019line) introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors.
+LaneATT(tabelini2021keep) effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module.
+CLRNet(zheng2022clrnet) employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features.
 This hierarchical fusion of semantic information significantly improves detection precision.
-FLAMNet[32] employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms.
+FLAMNet(ran2023flamnet) employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms.
 This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures.
 However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
-Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.[33]
-Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[33]
 2.4 Row-wise-based Lane Detection
 Row-wise methods predict lane positions for each image row.
-E2E-LMD[34] introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly.
-UFLD[35] utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes.
-CondLaneNet[36] utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts.
+E2E-LMD(yoo2020end) introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly.
+UFLD(qin2020ultra) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes.
+CondLaneNet(liu2021condlanenet) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts.
 The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
-3 Data Constrution
-Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models.
-In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions.
-Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging.
-Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability.
-However, there is currently no publicly available foggy weather lane detection dataset.
-To fill this gap, we created the FoggyLane dataset for lane detection in foggy weather.
-In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [8] and Tusimple [24] datasets to test the generalizability of our methods.
-Figure 4: The different road types of FoggyLane images examples.
-Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
-Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [37] 1.2k None 640 x 480 Tusimple [24] 6.4k None 1280 x 720 LLAMAS [38] 100k None 1276 x 717 Appolloscape [39] 110k None 3384 x 2710 BDD100k [40] 100k None 1280 x 720 CULane [8] 133k None 1640 x 590 CurveLanes [41] 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.423k 1640 x 590
-3.1 FoggyLane
-Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents.
-Current benchmark datasets, including CULane[8], Tusimple[24], and LLAMAS[38], lack foggy weather data.
-To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions.
-Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024.
-The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
-Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos[42], resulting in an additional 337 images.
-The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad.
-The dataset was partitioned into training, validation and testing sets, with stratification to maintain a consistent distribution of scene types across both subsets.
-Comparing FoggyLane with existing datasets, such as Caltech-Lanes[37] and Tusimple[24], which focus on single traffic scenarios, and LLAMAS[38], which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
-While ApolloScape [39] and BDD100k [40] include multiple traffic scenarios, they lack instance-level annotations.
-Similarly, although CULane [8] and CurveLanes [41] cover a wide range of traffic scenes, they do not include foggy conditions.
-Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced.
-To fill this gap, FoggyLane offers a thorough solution for lane detection in foggy conditions with instance-level annotations.
-Our dataset has two key advantages:
-  • 1)
-    Pioneering foggy lane detection dataset: FoggyLane is a dataset specifically designed to address foggy weather scenarios.
-It covers a range of challenging conditions, including foggy congested roads, foggy curves, and foggy night roads.
-This dataset is crucial for testing the robustness of ADAS under adverse conditions.
-  • 2)
-    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[43].
-These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
-Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
-Figure 6: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
-3.2 FoggyCULane and FoggyTusimple
-Currently, numerous datasets are available for lane detection.
-Among them, we select two widely adopted and representative benchmarks: CULane[8] and TuSimple[24].
-The CULane dataset [8] serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing.
-In contrast, the TuSimple dataset [24] primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
-Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
-The method for generating foggy images is inspired by the model [18], and the complete generation process is shown in Fig.
-5.
-To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions.
-It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one.
-The mathematical modeling equation defined by Koschmieder [44] is represented as:
-I(x)=J(x)t(x)+A(1t(x))I(x)=J(x)t(x)+A(1-t(x)) (1)
-where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image.
-$A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
-$t(x)=e^{-\beta d(x)}$ (2)
-where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter.
-$d(x)$ represents the distance between the object and the viewer.
-As this distance increases, the transmittance decreases, resulting in thicker fog.
-To calculate the corresponding $d(x)$, we use the S2R-DepthNet[45] that has good generalization performance across various applications.
-The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image.
-Then they suppress information irrelevant to depth and use structural information for depth prediction.
-The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details.
-The Depth-specific Attention (DSA) generates an attention map from the original image.
-Once processed through the DSA and STE Modules, as depicted in Fig.
-5, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map.
-The Depth Prediction (DP) module is then responsible for estimating the depth.
-Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
-According to the transmittance formula Eq.
-(2), we use the dark channel prior[46] to calculate the average atmospheric light value $A$.
-In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel.
-Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
-$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (3)
-where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$.
-In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$.
-After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$.
-The original images and the foggy images generated under different $\beta$ values are displayed in Figs.
-6 and 7.
+Figure 2: Overview of the proposed network.
+Figure 3: Backbone Network with Global Feature Fusion Module(GFFM).
 3 Methods
 3.1 Overall structure
-To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network.
+To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose PDT-Net, a robust Prior-Guided Dynamic Feature Fusion Transformer framework.
 The network works through several collaborative modules to tackle the challenges posed by foggy images, with the overall architecture shown in Fig.
-8.
+2.
 The input image first passes through the backbone network, where feature extraction occurs, followed by the Global Feature Fusion Module (GFFM).
@@
 The output is a feature map enriched with both local and global information, crucial for accurate lane detection.
-The multi-level feature map from the backbone is then passed into the Neck section of the network, which employs HRNetV2[47] for multi-scale feature fusion.
+The multi-level feature map from the backbone is then passed into the Neck section of the network, which employs HRNetV2(9052469) for multi-scale feature fusion.
 HRNetV2 excels in integrating high-resolution and low-resolution features from different stages of the backbone.
@@
 This approach eliminates the need for complex post-processing, significantly enhancing detection speed and accuracy.
-The output parameter map is then passed to the Kernel Feature Fusion Module (KFFM), which learns the relationships between different lane instances.
-The KFFM fuses these instance-specific features, improving lane localization accuracy, especially for complex or overlapping lane lines.
+The output parameter map is then passed to the Dynamic Feature Fusion Module (DFFM), which learns the relationships between different lane instances.
+The DFFM fuses these instance-specific features, improving lane localization accuracy, especially for complex or overlapping lane lines.
 This module ensures that the network can better identify lane instances and refine the localization of each lane.
-Finally, to compensate for the loss of edge information in foggy images, the network incorporates a Low-Level Edge Enhancement Module (LEEM).
+Finally, to compensate for the loss of edge information in foggy images, the network incorporates a Prior-Guided Edge Enhancement Module (PEM).
 This module enhances edge features using convolution operations, recovering fine details at lane boundaries.
 The enhanced edge features improve the network’s ability to detect lane boundaries with high precision, even under challenging foggy conditions.
-Figure 8: Overview of the proposed network.
 3.2 Global Feature Fusion Module
 As shown in Fig.
-9, to enhance global feature integration in the Swin Transformer backbone[48], we replace the final window-based attention mechanism with a standard multi-head attention mechanism (MHSA) from Vision Transformer (ViT)[49].
+3, to enhance global feature integration in the Swin Transformer backbone(liu2021swin), we replace the final window-based attention mechanism with a standard multi-head attention mechanism (MHSA) from Vision Transformer (ViT)(vit).
 While ViT excels at capturing global dependencies, it suffers from high computational cost, especially at high resolutions, making it unsuitable for real-time tasks such as lane detection.
@@
 In our approach, we address this issue by introducing an efficient solution: global attention is applied only after reducing the image resolution.
-This ensures that the computational cost of global attention is manageable, allowing us to retain the global attention mechanism of ViT while maintaining fast inference speed, which is crucial for real-time lane detection in foggy conditions.
+This ensures that the computational cost of global attention is manageable(AAP2), allowing us to retain the global attention mechanism of ViT while maintaining fast inference speed, which is crucial for real-time lane detection in foggy conditions.
+Figure 4: The structure of the Dynamic Feature Fusion Module(DFFM)
 In our design, this MHSA layer replaces the original window-based attention mechanism used in the Swin Transformer.
 By applying global attention after reducing the image resolution, we strike a balance between maintaining global context awareness and achieving high computational efficiency, which is crucial for real-time lane detection tasks.
-Figure 9: Backbone Network with Global Feature Fusion.
-4.3 Kernel Feature Fusion Module
-Existing lane instance segmentation methods[50, 36], typically use dynamic convolution to generate kernel parameters for each instance, which improves the network’s flexibility and adaptability.
+3.3 Dynamic Feature Fusion Module
+Existing lane instance segmentation methods(tian2020conditional; liu2021condlanenet), typically use dynamic convolution to generate kernel parameters for each instance, which improves the network’s flexibility and adaptability.
 However, these methods do not fully leverage the structural relationships between lane instances, leading to limited segmentation and detection accuracy.
 Traditional dynamic convolution methods only consider individual lane instances and overlook the potential relationships between adjacent instances, which reduces their performance when handling complex lane scenes.
-To address this limitation, we propose the Kernel Feature Fusion Module (KFFM), as shown in the bottom right of Fig.
-8, which optimizes kernel generation by integrating information from adjacent lane instances.
-Specifically, KFFM introduces structured feature fusion into dynamic convolution, allowing each lane instance’s kernel to depend not only on the features of the current lane but also on the information from nearby lane instances.
+To address this limitation, we propose the Dynamic Feature Fusion Module (DFFM), as shown in Fig.
+4, which optimizes kernel generation by integrating information from adjacent lane instances.
+Specifically, DFFM introduces structured feature fusion into dynamic convolution, allowing each lane instance’s kernel to depend not only on the features of the current lane but also on the information from nearby lane instances.
 This approach enables the network to adaptively learn the relationships between lane instances, improving both detection accuracy and segmentation performance.
@@
 The resulting fused kernel $\hat{\mathit{kernel}}_{i}$ is then used in the dynamic convolution operation for lane segmentation, leading to more accurate lane detection and segmentation.
-4.4 Low-level Edge Enhanced Module
-Images captured in foggy conditions differ from those captured in good lighting primarily due to blurriness and low contrast, which makes the edges of objects and scene details indistinct and difficult to detect.
-This degradation of edge information significantly impacts the algorithm’s ability to accurately understand and recognize the scene.
-Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial relationships about objects, which is critical for tasks such as object detection, segmentation, and tracking.
-We used the Canny edge detection algorithm with identical threshold settings to process both foggy and normal images, as shown in Fig.
-10.
-The results show that fog significantly degrades edge information, particularly lane line edges, which become nearly indiscernible.
-Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—accurate edge detection is essential for reliable lane detection.
-To address the challenges posed by foggy conditions, we propose the integration of global feature fusion within both the backbone network and detection head to enhance high-dimensional global features.
-Additionally, we incorporate multi-scale fusion to aggregate detailed features from low-resolution feature maps.
-However, this approach alone does not fully compensate for the loss of edge details in foggy images.
-To resolve this issue, we introduce the Low-level Edge Enhanced Module (LEEM) into the network.
-LEEM focuses on edge information, effectively recovering and refining the edges that are crucial for lane detection, thus improving both the accuracy and stability of lane detection in foggy environments.
+3.4 Prior-Guided Edge Enhancement Module
+In foggy conditions, images are typically characterized by reduced contrast and sharpness, making it difficult to detect key features such as edges5, which are critical for tasks like lane detection.
+In these scenarios, edge information becomes an essential prior knowledge that can guide the network in recovering vital details.
+Edge information not only defines object boundaries but also conveys structural and spatial relationships, crucial for understanding the scene.
+Recognizing this, we propose the Prior-Guided Edge Enhancement Module (PEM),as shown in Fig.
+6, which integrates prior knowledge of edge structures to recover lost edge information in foggy images, ultimately enhancing lane detection accuracy and robustness.
 Figure 5: Comparison of edge information between foggy and normal scenarios.
-LEEM works by adding an auxiliary path to the network that specifically learns edge features of the input image.
+PEM works by adding an auxiliary path to the network that specifically learns edge features of the input image.
 The process begins with the application of the Canny operator for edge detection, followed by the reconstruction of missing lane edges using lane line labels.
@@
 The mathematical representation of this process is as follows:
-$\hat{\mathit{E}}_{\mathit{predict}}=\mathit{Conv}_{2}\left(\mathit{ReLU}\left(\mathit{BN}(\mathit{Conv}_{1}(\mathit{X})))\right)\right)$ (6)
+$\hat{\mathit{E}}_{\mathit{predict}}=\mathit{Conv}_{2}\left(\mathit{ReLU}\left(\mathit{BN}(\mathit{Conv}_{1}(\mathit{X}))\right)\right)$ (3)
 where $\mathit{Conv}_{1}$ is a $3\times 3$ convolution designed to capture spatial edge features, $\mathit{BN}$ is the batch normalization layer to stabilize the learning process by normalizing the output, $\mathit{ReLU}$ is the activation function that provides non-linearity, $\mathit{Conv}_{2}$ is a $1\times 1$ convolution used to reduce the dimensionality and refine edge features.
@@
 where $e_{\mathit{gt}}^{i}$ represents the ground truth label for pixel $i$ (1 if it’s part of the edge, 0 otherwise), $p_{i}$ is the predicted probability that pixel $i$ is part of the edge.
-The LEEM module allows the network to effectively enhance critical edge details in lane detection, significantly improving both the accuracy and stability of lane detection in foggy conditions.
+The PEM module allows the network to effectively enhance critical edge details in lane detection, significantly improving both the accuracy and stability of lane detection in foggy conditions.
 This module not only recovers lost edge information in images but also integrates information from various sources, using the cross-entropy loss function to optimize the network.
-As a result, LEEM plays a key role in enhancing the robustness of lane detection systems in challenging environments.
-Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
-[…]
-  • 1.
-    Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
-* indicates the FPS after acceleration with TensorRT.
-Figure 11: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyLane testing set.
-The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
-Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
-[…]
-  • 1.
-    Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow and Night conditions, as well as in the overall (Total) performance.
+As a result, PEM plays a key role in enhancing the robustness of lane detection systems in challenging environments.
+Figure 6: The structure of the Prior-Guided Edge Enhancement Module(PEM)
 3.5 Loss functions
@@
 where $e_{gt}^{i}$ denotes the label of pixel $i$, and $p_{i}$ represents the predicted probability of the pixel being part of the edge.
+Figure 7: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(googlemap)
+Figure 8: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(googlemap)
 3.5.6 Total loss
@@
 4.1 Datasets
-In the Data Construction chapter, we introduce the FoggyLane dataset, which was constructed from real-world collections for foggy lane detection, along with the FoggyCULane and FoggyTusimple datasets that were synthesized through modeling foggy images.
-Our method is evaluated on these three datasets.
-FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
-FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach.
-It maintains the same number of images as CULane.
-The resolution of all images is also 1640x590 pixels.
-FoggyTusimple is synthesized from Tusimple through the foggy image modeling method.
-At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases.
-All visual data maintains a consistent resolution of 1280×720 pixels.
+Currently, there is no publicly available foggy weather lane detection dataset.
+To fill this gap, we created the FoggyLane dataset for lane detection in foggy weather.
+In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (pan2018spatial) and Tusimple (Tusimple) datasets to test the generalizability of our methods.
+Figure 9: The different road types of FoggyLane images examples.
+Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
+Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Caltech-Lanes) 1.2k None 640 x 480 Tusimple (Tusimple) 6.4k None 1280 x 720 LLAMAS (LLAMAS) 100k None 1276 x 717 Appolloscape (Apolloscape) 110k None 3384 x 2710 BDD100k (yu2020bdd100k) 100k None 1280 x 720 CULane (pan2018spatial) 133k None 1640 x 590 CurveLanes (xu2020curvelane) 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.423k 1640 x 590
+4.1.1 FoggyLane
+Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents.
+Current benchmark datasets, including CULane(pan2018spatial), Tusimple(Tusimple), and LLAMAS(LLAMAS), lack foggy weather data.
+To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions.
+Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024.
+The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
+Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube), resulting in an additional 337 images.
+The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad.
+The dataset was partitioned into training, validation and testing sets, with stratification to maintain a consistent distribution of scene types across both subsets.
+Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Caltech-Lanes) and Tusimple(Tusimple), which focus on single traffic scenarios, and LLAMAS(LLAMAS), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios.
+While ApolloScape (Apolloscape) and BDD100k (yu2020bdd100k) include multiple traffic scenarios, they lack instance-level annotations.
+Similarly, although CULane (pan2018spatial) and CurveLanes (xu2020curvelane) cover a wide range of traffic scenes, they do not include foggy conditions.
+Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced.
+To fill this gap, FoggyLane offers a thorough solution for lane detection in foggy conditions with instance-level annotations.
+Our dataset has two key advantages:
+  • 1)
+    Pioneering foggy lane detection dataset: FoggyLane is a dataset specifically designed to address foggy weather scenarios.
+It covers a range of challenging conditions, including foggy congested roads, foggy curves, and foggy night roads.
+This dataset is crucial for testing the robustness of ADAS under adverse conditions.
+  • 2)
+    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.(zhang2024robust).
+These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
+Figure 10: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
+Figure 11: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
+4.1.2 FoggyCULane and FoggyTusimple
+Currently, numerous datasets are available for lane detection.
+Among them, we select two widely adopted and representative benchmarks: CULane(pan2018spatial) and TuSimple(Tusimple).
+The CULane dataset (pan2018spatial) serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing.
+In contrast, the TuSimple dataset (Tusimple) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations.
+Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
+The method for generating foggy images is inspired by the model (nie2022sensor), and the complete generation process is shown in Fig.
+10.
+To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions.
+It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one.
+The mathematical modeling equation defined by Koschmieder (harald1924theorie) is represented as:
+I(x)=J(x)t(x)+A(1t(x))I(x)=J(x)t(x)+A(1-t(x)) (13)
+where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image.
+$A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
+$t(x)=e^{-\beta d(x)}$ (14)
+where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter.
+$d(x)$ represents the distance between the object and the viewer.
+As this distance increases, the transmittance decreases, resulting in thicker fog.
+To calculate the corresponding $d(x)$, we use the S2R-DepthNet(chen2021S2R) that has good generalization performance across various applications.
+The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image.
+Then they suppress information irrelevant to depth and use structural information for depth prediction.
+The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details.
+The Depth-specific Attention (DSA) generates an attention map from the original image.
+Once processed through the DSA and STE Modules, as depicted in Fig.
+10, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map.
+The Depth Prediction (DP) module is then responsible for estimating the depth.
+According to the transmittance formula Eq.
+(14), we use the dark channel prior(he2011darkchannel) to calculate the average atmospheric light value $A$.
+In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel.
+Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
+$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (15)
+where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$.
+In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$.
+After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$.
+The original images and the foggy images generated under different $\beta$ values are displayed in Figs.
+11.
+Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
+[…]
+  • Note: For a fairer comparison, we re-evaluated the FPS of the source code available detectors using one NVIDIA GeForce RTX 3090 GPU on the same machine.
+* indicates the FPS after acceleration with TensorRT.
 4.2 Evaluation Metric
@@
 However, due to its distinct characteristics, lane detection also requires specialized evaluation metrics.
-Adhering to SCNN’s [8] framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions.
+Adhering to SCNN’s (pan2018spatial) framework, we link labeled or predicted lane points to form line segments and extend them into 30-pixel-wide regions.
 Subsequently, we compute the Intersection over Union (IoU), and if the IoU value exceeds a predefined threshold, typically set at 0.5, the prediction is considered correct.
@@
 $Recall=\frac{TP}{TP+FN}$ (18)
-Figure 12: Visualization results of CLRNet[16], FLAMNet[32], PolarRCNN[51] and our method on FoggyCULane testing set.
+Figure 12: Visualization results of CLRNet(zheng2022clrnet), FLAMNet(ran2023flamnet), PolarRCNN(tits4) and our method on FoggyLane testing set.
+The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
+Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
+[…]
+  • Note: Our method attains state-of-the-art (SOTA) results in multiple scenarios including Normal, Crowded, Shadow, No-line, Arrow and Night conditions, as well as in the overall (Total) performance.
+Figure 13: Visualization results of CLRNet(zheng2022clrnet), FLAMNet(ran2023flamnet), PolarRCNN(tits4) and our method on FoggyCULane testing set.
 The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
@@
 For full network parameters and training details, please refer to our open-source code.
-To benchmark against existing methods[8, 35, 36, 25, 31, 28, 16, 32, 29, 51], we followed the training settings provided in the respective papers.
+To benchmark against existing methods(pan2018spatial; qin2020ultra; liu2021condlanenet; zheng2021resa; tabelini2021keep; liu2021end; zheng2022clrnet; ran2023flamnet; dai2024dbnet; tits4), we followed the training settings provided in the respective papers.
 For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results
@@
 4.5 Ablation Study
-We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone.
+We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, DFFM, and PEM modules, using Swin-t as the backbone.
 As shown in Tables 5 and 6, each module improves the model’s performance.
 The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information.
-Adding the KFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane.
-Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
+Adding the DFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane.
+Finally, the PEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
 Table 5: Ablation experiments of different components on the FoggyLane dataset.
-  Model GFFM KFFM LEEM F1-score Gflops(G)
+  Model GFFM DFFM PEM F1-score Gflops(G)
 Baseline 90.36 13.77
@@
 Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig.
-14.
+15.
 This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit.
@@
 Table 6: Ablation experiments of different components on the FoggyCULane dataset.
-  Model GFFM KFFM LEEM F1-score Gflops(G)
+  Model GFFM DFFM PEM F1-score Gflops(G)
 Baseline 77.85 13.77
@@
 […]
-Figure 14: A edge computing system built with NVIDIA Jetson AGX Orin.
-After experimental validation, it has been demonstrated that our method can run accurately on the vehicle’s edge computing platform with limited computational resources, achieving a performance of 38.4 FPS.
+Figure 15: Lane departure warning module within the active safety warning systems framework.
+The module detects unintentional lane departure using a forward-facing camera and edge computing system built with NVIDIA Jetson AGX Orin for real-time lane marking recognition.
+Upon detecting a departure trajectory (represented by yellow dotted lines), the system triggers multi-modal warnings including steering wheel vibration, seat vibration, and voice notification with corresponding warning icons on the dashboard monitor to proactively prevent accidents.
 5 Conclusion
-This study focuses on lane detection in foggy conditions, aiming to address the range of challenges posed by foggy environments, which is crucial for ensuring driving safety in fog and enhancing Advanced Driver Assistance System (ADAS) capabilities.
-To address the current lack of publicly available datasets for lane detection in foggy conditions, we constructed a real-world foggy lane dataset, FoggyLane, and synthesized two additional foggy datasets, FoggyCULane and FoggyTusimple, based on the popular CULane and Tusimple datasets.
-In response to the challenges of lane detection in foggy scenarios, we propose an end-to-end fog-enhanced lane detection method.
-First, this method incorporates a Global Feature Fusion Module in the backbone network to address the absence of global spatial information in foggy scenes.
-Second, we employ a dynamic convolution-based detection head for instance-sensitive lane segmentation, integrating a Kernel Feature Fusion Module within the detection head to automatically capture associations between lane instances.
-Additionally, a Low-Level Edge Enhancement Module was added to increase the network’s focus on edge information, mitigating the loss of edge details and improving lane detection accuracy under foggy conditions.
-Comparative experiments on the FoggyLane, FoggyCULane, and FoggyTusimple datasets with other advanced lane detection methods demonstrate that our approach achieves superior performance in both clear and foggy conditions, significantly improving lane detection accuracy.
-Furthermore, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, further validating its practical applicability in real-world scenarios.
+This study addresses the critical challenge of lane detection in foggy conditions, which is essential for improving road safety and enhancing Advanced Driver Assistance Systems (ADAS).
+To tackle the issue of insufficient publicly available datasets for foggy lane detection, we constructed the FoggyLane dataset, a real-world dataset specifically targeting lane detection in foggy environments, and synthesized two additional datasets, FoggyCULane and FoggyTusimple, to fill the gap in fog-specific data.
+In response to the difficulties posed by foggy scenarios, we propose a robust Prior-Guided Dynamic Feature Fusion Transformer framework for real-time lane detection.
+This method incorporates a Global Feature Fusion Module to capture the relationship between local and global features in foggy images, a Dynamic Feature Fusion Module to model the structural and positional relationships of lane instances, and a Prior-Guided Edge Enhancement Module to recover lost edge details.
+Extensive experiments show that our approach significantly improves lane detection accuracy compared to other advanced methods, both in clear and foggy conditions, with F1-scores of 95.04%, 79.85%, and 96.95% on the FoggyLane, FoggyCULane, and FoggyTusimple datasets, respectively.
+Moreover, with TensorRT acceleration, our method achieves an inference speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its practical applicability in real-world scenarios.
+By improving the precision of lane detection, our framework contributes to active safety warning systems, helping prevent accidents in foggy conditions and enhancing road safety, particularly in challenging driving environments.
 Acknowledgement
@@
 2023B1212060029).We would like to release our source code and dataset.
+6 Data Availability
+Data will be made available on request.
+7 Declaration of Competing Interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
+8 Acknowledgement
+This project is jointly supported by National Natural Science Foundation of China (Nos.52172350, 51775565), Guangdong Basic and Applied Research Foundation (Nos.2
+021B1515120032, 2022B1515120072), Guangzhou Science and Technology Plan Project (Nos.2024B01W0079,20
+2206030005), Science and Technology Planning Project of Guangdong Province (No.2023B1212060029).
+[…]